### v1

# A unified approach to hypergeometric class functions

Abstract
Hypergeometric class equations are given by second order differential operators in one variable whose coefficient at the second derivative is a polynomial of degree $\leq 2$, at the first derivative of degree $\leq 1$ and the free term is a number. Their solutions, called hypergeometric class functions, include the Gauss hypergeometric function and its various limiting cases. The paper presents a unified approach to these functions. The main structure behind this approach is a family of complex 4-dimensional Lie algebras, originally due to Willard Miller. Hypergeometric class functions can be interpreted as eigenfunctions of the quadratic Casimir operator in a representation of Miller’s Lie algebra given by differential operators in three complex variables. One obtains a unified treatment of various properties of hypergeometric class functions such as recurrence relations, discrete symmetries, power series expansions, integral representations, generating functions and orthogonality of polynomial solutions.
1 Introduction
This paper is devoted to the family of equations of the form
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta\right)f(z)=0,$ (1.1)
where $\sigma(z),\ \tau(z),\ \eta$ are polynomials with
${\deg}\sigma\leq 2,\ {\deg}\tau\leq 1,\ \ \deg\eta=0.$ (1.2)
Their solutions include some of the most useful special functions with many applications in physics and mathematics.
In the literature one can find several names for this family. In this paper, we will use the name hypergeometric class equations for equations of the form (1.1) satisfying (1.2).111 In the book by Nikiforov-Uvarov [NU], and also in [De1, De2], this family is called equations of the hypergeometric type. However, Slavianov-Lay’s book [SL] suggests to use the term “type” for smaller families, such as the families (1)–(5) listed below. [DIL] calls this family grounded hypergeometric class. The name hypergeometric class, following the suggestions of Slavianov-Lay, is reserved in [DIL] for a somewhat wider family, where $\eta(z)$ is allowed to be a rational function such that $\eta\sigma$ is a polynomial of degree $\leq 2$. Anyway, in this paper for simplicity we drop the adjective “grounded”.
Let us start with a short review of basic nontrivial types of hypergeometric class equations. We will always assume that $\sigma\neq 0$. Every class will be simplified by dividing by a constant and an affine change of the complex variable $z$.
  • (1)
    The ${}_{2}F_{1}$ or Gauss hypergeometric equation
    $\left(z(1-z)\partial_{z}^{2}+(c-(a+b+1)z)\partial_{z}-ab\right)f(z)=0.$
  • (2)
    The ${}_{2}F_{0}$ equation
    $\left(z^{2}\partial_{z}^{2}+(-1+(1+a+b)z)\partial_{z}+ab\right)f(z)=0.$
  • (3)
    The ${}_{1}F_{1}$ or Kummer’s confluent equation
    $(z\partial_{z}^{2}+(c-z)\partial_{z}-a)f(z)=0.$
  • (4)
    The ${}_{0}F_{1}$ equation, closely related to the Bessel equation,
    $(z\partial_{z}^{2}+c\partial_{z}-1)f(z)=0.$
  • (5)
    The Hermite equation
    $(\partial_{z}^{2}-2z\partial_{z}-2a)f(z)=0.$
In our work we collect various results about hypergeometric class equations that can be stated and proven in a unified way, with as few restrictions on $\sigma$, $\tau$, $\eta$ as possible. We believe that such an approach has a considerable pedagogical and theoretical value. The pedagogical advantage of the unified approach is obvious: it reduces the need for repetitive arguments. From the theoretical point of view, in this way we easily see the coalescence of various types. The properties of hypergeometric class equations that we describe depend analytically on the the coefficients $\sigma,\tau,\eta$. These properties include a pair of recurrence relations, a discrete symmetry, integral representations, power expansions around singular points, generating functions, the Rodriguez formulas for polynomials and their orthogonality.
The unified approach has its limitations. There are some properties that are not easy to formulate in a unified way. For instance, only one pair of recurrence relations depends analytically on the coefficients. If we restrict ourselves to specific types, then we often find a bigger number of recurrence relations (eg. at least 12 in the case of the Gauss hypergeometric equation). Another family of important properties not included in our presentation are quadratic identities, which link various types of hypergeometric class functions. Thus this work should be compared with other studies of hypergeometric class equations, such as [De1], where specific types are described one by one.
[…]
where $\alpha,\beta$ are complex parameters.
Let $\sigma$ be a polynomial of degree $\leq 2$ (as in (1.1)) and $\kappa$ a polynomial of degree $\leq 1$. Miller’s Lie algebra with $\alpha=\frac{\sigma^{\prime\prime}}{2}$ and $\beta=\kappa^{\prime}$ can be represented by the following 1st order differential operators acting on ${\mathbb{C}}^{3}$:
$\begin{array}[]{l}N:=t\partial_{t}-s\partial_{s},\\[5.69054pt] A_{+}:=t\partial_{z}+\sigma^{\prime}(z)\partial_{s},\\[5.69054pt] A_{-}:=s\partial_{z}+\sigma^{\prime}(z)\partial_{t}+\frac{\kappa(z)}{t}.\end{array}$ (1.4)
The operators (1.4) can be restricted to functions on the quadric
$\sigma(z)-ts=0.$ (1.5)
Miller’s Lie algebra $m_{\alpha,\beta}$ commutes with
${\mathcal{C}}_{\alpha,\beta}:=\frac{1}{2}(A_{-}A_{+}+A_{+}A_{-})+\alpha N^{2}+\beta N,$ (1.6)
which will be called the Casimir operator of Miller’s Lie algebra. In the representation (1.4), restricted to the quadric 1.5 and the eigenspace $N=0$, we obtain a 2nd order differential operator in $z$, denoted ${\mathcal{C}}(\sigma,\kappa)$. Its eigenvalue equation
$\big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\big{)}f=0$ (1.7)
is precisely the hypergeometric class equation (1.1), where
$\tau(z)=\kappa(z)+\sigma^{\prime}(z),\quad\eta=\frac{\kappa^{\prime}}{2}+\omega.$ (1.8)
The operators $A_{+}$ and $A_{-}$ can be treated as “root operators” wrt the “Cartan algebra” spanned by $N$. They lead to recurrence relations for ladders of solutions. The involution of the quadric (1.5) $(t,s,z)\mapsto(s,t,z)$ generates a “Weyl symmetry” of Miller’s Lie algebra. It leads to a discrete symmetry of hypergeometric class equations.
We devote Section 3 to the properties of hypergeometric class equations and functions which follow directly from the properties of Miller’s Lie algebra: the basic recurrence relations and the basic symmetry. Thanks to the recurrence relations, once we know a solution of a certain hypergeometric class equation, we know its solutions for a whole ladder of parameters labelled by $n\in{\mathbb{Z}}$.
Certain ladders are special. One of them contains solutions which can be expressed in terms of elementary functions. We call them Chebyshev solutions, because it contains the well-known Chebyshev polynomials. It is described in Subsection 3.6.
Another ladder of solutions consists of polynomials. In the convention that we adopted, this ladder is descending—its elements are zero for $n>0$. We devote to this ladder the whole Section 7. This family includes the famous classical orthogonal polynomials. The basic symmetry produces from the polynomial ladder an ascending ladder, which needs no separate discussion.
For the Hermite equation the polynomial $\sigma$ does not have a zero. Other types, that is equations reducible to the ${}_{2}F_{1}$, ${}_{1}F_{1}$, ${}_{2}F_{0}$ and ${}_{0}F_{1}$ types, $\sigma$ has at least one zero. We devote Section 4 to these equations, where without loss of generality we assum $\sigma(0)=0$. We can then write a formal power series $F(\sigma,\kappa,\omega;z)$ which solves the equation
$\big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\big{)}F=0,$ (1.9)
normalized to be $1$ at zero. This power series is convergent if $\sigma^{\prime}(0)\neq 0$, that is for ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$. If $\sigma^{\prime}(0)=0$, that is for ${}_{2}F_{0}$, it is divergent, but it is asymptotic to a well-defined solution. Thus we obtain a unification of 4 types of hypergeometric functions in a single function $F(\sigma,\kappa,\omega;z)$, which we call the unified hypergeometric function. It depends meromorphically on $5$ complex parameters (two parameters for $\sigma$, two for $\kappa$ and one for $\omega$).
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, it is natural to assume $\sigma^{\prime}(0)=1$ and to normalize the unified hypergeometric function by dividing it by $\Gamma(m+1)$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. This is sometimes called Olver’s normalization. Thus we obtain the Olver’s normalized unified hypergeometric function $\mathbf{F}(\sigma,\kappa,\omega;z)$, which depends analytically on $4$ complex parameters (one parameter for $\sigma$, two for $\kappa$ and one for $\omega$).
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$ the hypergeometric class equation has an additional discrete symmetry, which involves the gauge transformation $x^{m}\cdot x^{-m}$. We call it the power symmetry.
If $\sigma(0)=0$, and $\sigma^{\prime\prime}\neq 0$ or $\kappa^{\prime}\neq 0$, that is in the cases of the ${}_{2}F_{1}$, ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, we have yet another discrete symmetry, which we call the inversion symmetry. For the ${}_{2}F_{1}$ equation it involves the change of the variable $z\mapsto\frac{1}{z}$ and the transformed equation is still of the form ${}_{2}F_{1}$. In the confluent cases it interchanges the ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, and it involves $z\to-\frac{1}{z}$. It is however not defined for the ${}_{0}F_{1}$ equation.
In Section 5 we describe a unified approach to representations of hypergeometric class functions in terms of integrals of elementary functions. The ${}_{2}F_{1}$, ${}_{1}F_{1}$, ${}_{2}F_{0}$ and Hermite functions can be represented as Euler integrals, that is,
$\int_{\gamma}p(s)(s-z)^{-n-1}{\rm d}s,$ (1.10)
where $p$ is an elementary function, $n\in{\mathbb{C}}$ and $\gamma$ is an appropriate contour. ${}_{1}F_{1}$, ${}_{0}F_{1}$ and Hermite functions possess representations in the form of the Laplace integral
$\int_{\gamma}q(s){\rm e}^{zs}{\rm d}s,$ (1.11)
where $q$ is an elementary function and $\gamma$ is an appropriate contour.
As we mentioned above, some ladders of solutions of hypergeometric class equations consist of polynomials. These polynomials include the famous classical orthogonal polynomials (sometimes called the very classical orthogonal polynomials). They allow for a very elegant unified treatment, which includes not only recurrence relations, but also generating functions (the famous Rodrigues formula) and the orthogonality relations. All of this is briefly described in Section 7.
As we mentioned above, if $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, that is in the cases equivalent to types ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$, we have the power symmetry. Because of that, these equations have two linearly independent solutions with a distinct behavior at zero: the unified hypergeometric function, which is analytic at zero, and another solution behaving as $\sim z^{m}$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. The linear independence breaks down for $m\in{\mathbb{Z}}$, when both solutions are proportional to one another. This case is called degenerate and is discussed in Section 8. The unified hypergeometric function in this case has an additional integral representation and an elegant generating function.
The present work should be used only as an “invitation” to hypergeometric class functions. It leaves out many of their properties, which are difficult to describe in a unified way. For instance, as we mentioned above, various types of hypergeometric class functions possess additional recurrence relations and additional symmetries. However only those derived directly from Miller’s Lie algebra can be described in a fully unified fashion.
Almost all our discussion is algebraic, without functional analysis. We are aware that it is natural to view hypergeometric class operators as closed (or self-adjoint) Sturm-Liouville operators on appropriate weighted $L^{2}$ spaces. It would be also interesting to consider representations of Miller’s Lie algebra in Hilbert spaces. This would, however, require breaking our discussion into separate types. We use some elements of functional analysis only in Section 7. about classical orthogonal polynomials. We show how to view them as eigenfunctions of certain self-adjoint Sturm-Liouville operators on weighted $L^{2}$ spaces with appropriate boundary conditions—this can be done in a rather unified fashion.
The literature on hypergeometric class equations is very large. Usually, each type is considered separately, without an attempt of a unified treatment. Let us list some of the more famous treatises about these equations: [EMOT, Ho, MOS, MF, R, Wa, WW]
Under the name “equations of the hypergeometric type” they were considerd in a unified way in the book by Nikiforov–Uvarov [NU]. This book was one of the two main inspirations for our article. A part of the material of our work is adapted from [NU], notably the material of Subsections 3.5, 5.1, 7.2, 7.6.
We use also ideas of Miller, notably in Section 2. As we mentioned above, Miller’s Lie algebra was defined in [M1]. Miller was an early champion of the use of Lie algebras in the theory of special functions. To my knowledge, he was the first to note that various types of hypergeometric class admit a larger symmetry algebra. This topic was further developed in [De2].
Let us also mention earlier works that contained elements of a unified theory of hypergeometric class equations—[IH] and [Tr].
More complete analysis of various types hypergeometric class equations, including the Lie-algebraic interpretation of their recurrence relations and discrete symmetries, can be found in the literature, notably in the works of Miller, and also in [De1, De2].
The family of hypergeometric class polynomials that form an orthogonal basis in an appropriate weighted Hilbert space consists essentially of Jacobi, Laguerre and Hermite polynomials, often called classical orthogonal polynomials It has an especially large literature, e.g. [NU, R]. In the more recent literature the name “classical orthogonal polynomials” is sometimes given to a broader family, given by the so-called Askey scheme. Some authors proposed to use the name very classical orthogonal polynomials for the family consisting of Jacobi, Laguerre and Hermite polynomials.
2 Miller’s Lie algebra
In this section we introduce a certain two-parameter family of 4-dimensional Lie algebras, which to our knowledge was first introduced by Willard Miller. We will call it Miller’s Lie algebra and denote by $m_{\alpha,\beta}$ with $\alpha,\beta\in{\mathbb{C}}$. We will also describe its Casimir operator ${\mathcal{C}}_{\alpha,\beta}$, that is, a quadratic expression in elements of $m_{\alpha,\beta}$, which in any representation commutes with the whole Lie algebra. In other words, ${\mathcal{C}}_{\alpha,\beta}$ belongs to the center of the enveloping algebra of $m_{\alpha,\beta}$. We will also describe a family of representations of $m_{\alpha,\beta}$ by 1st order differential operators on certain 2nd degree surfaces in ${\mathbb{C}}^{3}$.
In appropriate coordinates the eigenvalue equation for the Casimir operator equation will have the form of a hypergeometric class equation. A number of properties of hypergeometric class equations will have a simple interpretation in terms of properties of Miller’s Lie algebra. They include the basic symmetry, basic factorizations and basic recurrence relations. They will be described in the next section.
2.1 Three low-dimensional Lie algebras
Let us first introduce three low-dimensional complex Lie algebras.
  1. 1.
    $sl(2,{\mathbb{C}})$. It is spanned by $N,A_{+},A_{-}$ satisfying the commutation relations
    $\begin{array}[]{l}[N,A_{\pm}]=\pm A_{\pm},\\[5.69054pt] [A_{+},A_{-}]=2N.\end{array}$ (2.1)
    The operator
    ${\mathcal{C}}:=$ $A_{-}A_{+}+N(N+1)$ (2.2)
    $=$ $A_{+}A_{-}+N(N-1)$ (2.3)
    $=$ $\frac{1}{2}(A_{-}A_{+}+A_{+}A_{-})+N^{2}$ (2.4)
    commutes with all elements of $sl(2,{\mathbb{C}})$. It is often called the Casimir (operator).
    Here is a typical representation of $sl(2,{\mathbb{C}})$:
    $N$ $:=x\partial_{x}-y\partial_{y},$ (2.5)
    $A_{+}$ $:=x\partial_{y},\quad A_{-}:=y\partial_{x}.$ (2.6)
  2. 2.
[…]
    The following operator commutes with $osc({\mathbb{C}})$, and will be called the Casimir of $osc({\mathbb{C}})$:
    ${\mathcal{C}}:=$ $A_{-}A_{+}+N+\tfrac{1}{2}$ (2.8)
    $=$ $A_{+}A_{-}+N-\tfrac{1}{2}$ (2.9)
    $=$ $\frac{1}{2}(A_{-}A_{+}+A_{+}A_{-})+N.$ (2.10)
    Here is a typical representation of $osc({\mathbb{C}})$:
    $N$ $:=\frac{1}{2}(-\partial_{x}^{2}+x^{2}),$ (2.11)
    $A_{\pm}$ $:=\frac{1}{\sqrt{2}}(x\mp\partial_{x}).$ (2.12)
    Note that $N$ is the quantum harmonic oscillator, which justifies the name of this Lie algebra.
  3. 3.
    The Lie algebra of Euclidean movements of the plane, denoted ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})$. It is spanned by $N,A_{+},A_{-}$ satisfying the commutation relations
    $\begin{array}[]{l}[N,A_{\pm}]=\pm A_{\pm},\\[5.69054pt] [A_{+},A_{-}]=0.\end{array}$ (2.13)
The following operator commutes with ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})$, and will be called the Casimir of ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})$:
${\mathcal{C}}:=$ $A_{-}A_{+}$ (2.14)
Here is a typical representation of ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})$:
$N$ $:=x\partial_{y}-y\partial_{x},$ (2.15)
$A_{\pm}$ $:=\partial_{x}\pm{\rm i}\partial_{y}.$ (2.16)
Thus $N$ is the generator of rotations of the plane and $A_{\pm}$ generate translations.
2.2 The family of Lie algebras introduced by W. Miller
The three Lie algebras introduced in the previous subsection can be joined in a single family depending on two complex parameters $\alpha,\beta$. This family will be denoted by $m_{\alpha,\beta}$. It was introduced by Willard Miller in [M1]. We will call it Miller’s Lie algebra.
[…]
It is easy to see that
$m_{\alpha,\beta}$ $\simeq sl(2,{\mathbb{C}})\oplus{\mathbb{C}},$ $\quad\alpha\neq 0;$ (2.18)
$m_{\alpha,\beta}$ $\simeq osc({\mathbb{C}}),$ $\quad\alpha=0,\quad\beta\neq 0;$ (2.19)
$m_{\alpha,\beta}$ $\simeq{\mathbb{C}}^{2}{\rtimes}so(2,{\mathbb{C}})\oplus{\mathbb{C}},$ $\quad\alpha=0,\quad\beta=0.$ (2.20)
Define the linear map $\tau:m_{\alpha,\beta}\to m_{\alpha,-\beta}$ by
[…]
Then $\tau$ is an isomorphism. The identity automorphism together with $\tau$ form a group isomorphic to ${\mathbb{Z}}_{2}$, which will be called the Weyl group of $m_{\alpha,\beta}$.
Similarly, for $n\in{\mathbb{C}}$, the linear map $\pi_{n}:m_{\alpha,\beta}\to m_{\alpha,\beta-2n\alpha}$ given by
[…]
is an isomorphism.
2.3 Casimir
Consider a representation of $m_{\alpha,\beta}$ on a vector space ${\cal Z}$. Following Miller [M1], introduce the following operator ${\mathcal{C}}$, called the Casimir of $m_{\alpha,\beta}$:
${\mathcal{C}}_{\alpha,\beta}={}{\mathcal{C}}:=$ $A_{-}A_{+}+\alpha N(N+1)+\beta(N+\tfrac{1}{2})$ (2.23)
$=$ $A_{+}A_{-}+\alpha N(N-1)+\beta(N-\tfrac{1}{2})$ (2.24)
$=$ $\frac{1}{2}(A_{-}A_{+}+A_{+}A_{-})+\alpha N^{2}+\beta N.$ (2.25)
As Miller noted, ${\mathcal{C}}$ commutes with the whole Lie algebra:
$N{\mathcal{C}}$ $={\mathcal{C}}N,\qquad A_{\pm}{\mathcal{C}}={\mathcal{C}}A_{\pm}.$ (2.26)
Extend the isomorphisms $\tau$ and $\pi_{n}$ defined in (2.21) and (2.22) to the algebra of operators on ${\cal Z}$. Then
$\tau({\mathcal{C}}_{\alpha,\beta})$ $={\mathcal{C}}_{\alpha,-\beta},$ (2.27)
[…]
2.4 Ladders
By a two-sided ladder we mean a subset of ${\mathbb{C}}$ of the form $n_{0}+{\mathbb{Z}}$, where $n_{0}\in{\mathbb{C}}$. A subset of the form $n_{0}+{\mathbb{N}}_{0}$ or $n_{0}-{\mathbb{N}}_{0}$ will be called a one-sided ladder (ascending or descending). A subset of the form $\{n_{0},n_{0}+1,\dots,n_{0}+n\}$ for some $n\in{\mathbb{N}}_{0}$ will be called a finite ladder.
It is easy to see that if ${\cal Z}$ is irreducible and $N$ possesses an eigenvalue, then the spectrum of $N$ is a ladder. In fact this follows from
$Nv=nv\quad\Rightarrow\quad NA_{\pm}v=(n\pm 1)A_{\pm}v.$ (2.29)
An irreducible representation with a one-sided or finite ladder with the lowest, resp. highest element 00 will be called a lowest, resp. heighest weight representation. It follows from (2.23), resp. (2.24) that
[…]
For $n\in{\mathbb{C}}$ we define
${\cal Z}^{n}:=\{v\in{\cal Z}\ :\quad\quad Nv=nv\}.$ (2.32)
Let ${\mathcal{C}}^{n}$, resp. $A_{\pm}^{n\pm\frac{1}{2}}$, be the operator ${\mathcal{C}}$, resp. $A_{\pm}$, restricted to ${\cal Z}^{n}$. (2.26) can be rewritten as
$A_{\pm}^{n\pm\frac{1}{2}}{\mathcal{C}}^{n}={\mathcal{C}}^{n}A_{\pm}^{n\pm\frac{1}{2}}.$ (2.33)
2.5 Representation by 1st order differential operators
Let $\sigma$ be a polynomial of degree $\leq 2$ and $\kappa$ of degree $\leq 1$. Consider ${\mathbb{C}}^{3}$ with elements denoted by $(t,s,z)$. Define the operators
$\begin{array}[]{l}N:=t\partial_{t}-s\partial_{s},\\[5.69054pt] A_{+}:=t\partial_{z}+\sigma^{\prime}(z)\partial_{s},\\[5.69054pt] A_{-}:=s\partial_{z}+\sigma^{\prime}(z)\partial_{t}+\frac{\kappa(z)}{t}.\end{array}$ (2.34)
[…]
$\alpha=\frac{\sigma^{\prime\prime}}{2},\quad\beta=\kappa^{\prime}.$ (2.36)
The operators (2.35) commute with the multiplication by $\sigma(z)-ts$. Therefore, we can restrict them to analytic functions on $\Omega$, the universal cover of the manifold
$\big{\{}(s,t,z)\in{\mathbb{C}}^{3}\ :\ \sigma(z)-ts=0,\ t\neq 0,\ s\neq 0\big{\}}.$ (2.37)
Let ${\cal A}(\Omega)$ denote the space of analytic functions on $\Omega$. The Lie algebra $m_{\alpha,\beta}$ represented by (2.35) acting on ${\cal A}(\Omega)$ will be denoted $m(\sigma,\kappa)$.
By (2.25), the Casimir of $m(\sigma,\kappa)$ is following differential operator on ${\cal A}(\Omega)$:
$\begin{array}[]{rl}{\mathcal{C}}=&st\partial_{z}^{2}+\sigma^{\prime}(z)(s\partial_{s}+t\partial_{t}+1)\partial_{z}+\kappa(z)\partial_{z}\\[5.69054pt] &+\frac{\sigma^{\prime\prime}}{2}\left((t\partial_{t})^{2}+(s\partial_{s})^{2}-2t\partial_{t}s\partial_{s}+t\partial_{t}+s\partial_{s}\right)\\[8.53581pt] &+\frac{(\sigma^{\prime}(z))^{2}}{ts}t\partial_{t}s\partial_{s}+\frac{\kappa(z)\sigma^{\prime}(z)}{ts}s\partial_{s}+\kappa^{\prime}(t\partial_{t}-s\partial_{s}+\frac{1}{2}).\end{array}$ (2.38)
Clearly, ${\mathcal{C}}$ commutes with $m(\sigma,\kappa)$.
Introduce new coordinates on ${\mathbb{C}}^{3}$:
$v=\sigma(z)-ts,\ \ \ \ w=t,\ \ \ \ \check{z}=z.$ (2.39)
The inverse transformation is
$\begin{array}[]{l}t=w,\ \ \ \ \ s=\frac{\sigma(z)-v}{w},\ \ \ \ \ z=\check{z}.\end{array}$
We have
$\begin{array}[]{l}\partial_{z}=\partial_{\check{z}}+\sigma^{\prime}(z)\partial_{v},\ \ \ \ \ \partial_{t}=\partial_{w}-s\partial_{v},\ \ \ \ \ \partial_{s}=-t\partial_{v}\end{array}$
Clearly, (2.37) is $v=0$. Therefore the operators (2.34) in new coordinates, after restricting to the surface (2.37) (and dropping “checks”), are
$\begin{array}[]{l}{}N=w\partial_{w},\\[5.69054pt] {}A_{+}=w\partial_{z},\\[5.69054pt] {}A_{-}=\frac{1}{w}\left(\sigma(z)\partial_{z}+\sigma^{\prime}(z)w\partial_{w}+\kappa(z)\right).\end{array}$ (2.40)
The Casimir operator is
$\begin{array}[]{ll}{}{\cal C}&={}A_{-}{}A_{+}+\frac{\sigma^{\prime\prime}}{2}{}N({}N+1)+\kappa^{\prime}{}(N+\frac{1}{2})\\[5.69054pt] &=\sigma(z)\partial_{z}^{2}+(\kappa(z)+\sigma^{\prime}(z)(w\partial_{w}+1))\partial_{z}+\frac{\sigma^{\prime\prime}}{2}w\partial_{w}(w\partial_{w}+1)+\kappa^{\prime}(w\partial_{w}+\frac{1}{2}).\end{array}$
Consistently with (2.32), for $n\in{\mathbb{C}}$ let
${\cal Z}^{n}:=\{g\in{\cal A}(\Omega)\ :\ \quad Ng=ng\}.$
As in the previous subsection, ${\mathcal{C}}^{n}$, resp. $A_{\pm}^{n\pm\frac{1}{2}}$ denote the operator ${\mathcal{C}}$, resp. $A_{\pm}$ restricted to ${\cal Z}^{n}$. Clearly, in the coordinates (2.39)
${\cal Z}^{n}=\{w^{n}F(z)\ :\quad F\in{\cal A}(\Theta)\},$
[…]
We have
${\cal C}^{n}+\omega:=\sigma(z)\partial_{z}^{2}+(\kappa(z)+\sigma^{\prime}(z)(n+1))\partial_{z}+\frac{\sigma^{\prime\prime}}{2}n(n+1)+\kappa^{\prime}(n+\tfrac{1}{2})+\omega.$ (2.41)
$A_{+}^{n+\frac{1}{2}}:=\partial_{z},\ \ \ \ A_{-}^{n-\frac{1}{2}}:=\sigma(z)\partial_{z}+\kappa(z)+n\sigma^{\prime}(z).$
2.6 Implementation of isomorphisms
Let us go back to ${\mathbb{C}}^{3}$. The implementation of the isomorphism $\pi_{n}$ defined in (2.22) is simple:
$\pi_{n}(\cdot)=t^{-n}\cdot t^{n}.$ (2.42)
Let us now describe an implementation of the isomorphism $\tau$. Let $\rho(z)$ solve
$(\sigma(z)\partial_{z}-\kappa(z))\rho(z)=0,$ (2.43)
(which defines $\rho(z)$ up to a coefficient). We introduce the following transformation on ${\cal A}(\Omega)$:
$Tg(t,s,z):=\rho(z)g(s,t,z).$
Theorem 2.1
$TNT^{-1}=-N,$ (2.44)
$TA_{+}T^{-1}=A_{-},$ (2.45)
$TA_{-}T^{-1}=A_{+}.$ (2.46)
$T{\cal C}T^{-1}={\cal C}.$ (2.47)
where on the left we have the operators from $m(\sigma,\kappa)$ and on the right from $m(\sigma,-\kappa)$.
Proof.  (2.44) is immediate.
Consider $g(t,s,z)\in{\cal A}(\Omega)$. Then
$\begin{array}[]{rl}T^{-1}g(t,s,z)&=\rho^{-1}(z)g(t,s,z),\\[8.53581pt] A_{+}T^{-1}g(t,s,z)&=(t\rho^{-1}(z)\partial_{z}-\frac{\kappa(z)}{\sigma(z)}t\rho^{-1}(z)+\sigma^{\prime}(z)\rho^{-1}(z)\partial_{s})g(s,t,z)\\[8.53581pt] &=(t\rho^{-1}(z)\partial_{z}-\frac{\kappa(z)}{s}\rho^{-1}(z)+\sigma^{\prime}(z)\rho^{-1}(z)\partial_{s})g(s,t,z)\\[8.53581pt] TA_{+}T^{-1}g(t,s,z)&=(s\partial_{z}-\frac{\kappa(z)}{t}+\sigma^{\prime}(z)\partial_{t})g(t,s,z).\end{array}$
This proves (2.45). The proof of (2.46) is similar.
To show (2.47) we use the second line of (2.38), (2.44), (2.45) and (2.46). $\Box$
In the coordinates (2.39) the symmetry equals
$\begin{array}[]{l}T\check{g}(w,z)=\rho(z)\check{g}(\frac{\sigma(z)}{w},z)\end{array}.$
Let $T^{n}$ be $T$ restricted to ${\cal Z}^{n}$. Clearly,
${\mathcal{C}}^{n}T^{n}={\mathcal{C}}^{-n}T^{-n}.$ (2.48)
We easily see that in the coordinates (2.39)
$T^{n}F(z):=\rho(z)\sigma^{n}(z)F(z).$
3 Basic properties of hypergeometric class equations
In this section we introduce hypergeometric class equations. They will be presented as the eigenvalue equation of a certain second order differential operator ${\mathcal{C}}(\sigma,\kappa)$. We discuss a number of properties of these equations that can be described in a unified way: the basic symmetry, basic factorization and basic recurrence relations.
The operator ${\mathcal{C}}(\sigma,\kappa)$ is essentially the Casimir of Miller’s Lie algebra at $N=0$, and all its properties discussed in this section follow from the properties of Miller’s Lie algebra analyzed in the previous section. However, this section can be read independently.
3.1 Remark about notation
Let $a,b,c$ be certain complex functions. Instead of saying that we consider the equation
$\Big{(}a(z)\partial_{z}^{2}+b(z)\partial_{z}+c(z)\Big{)}f(z)=0,$ (3.1)
we will usually say that the equation is given by the operator
${\cal A}:=a(z)\partial_{z}^{2}+b(z)\partial_{z}+c(z)$ (3.2)
Instead of ${\cal A}$, we will sometimes use the notation ${\cal A}(z,\partial_{z})$ to indicate the variable that is used in the given operator. This is useful when we consider a change of variables.
3.2 Parametrization of hypergeometric class operators
As was described in the introduction, the main topic of the paper are equations given by operators of the form
$\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta,$ (3.3)
where $\sigma,\tau,\eta$ satisfy the conditions (1.2). However, the parametrization of these equations with $\sigma,\tau,\eta$ is not always convenient. In this subsection we introduce another parametrization, which is usually better.
Let $\sigma(z),\ \kappa(z),\omega$ be polynomials with
${\deg}\sigma\leq 2,\ {\deg}\kappa\leq 1,\ \ \deg\omega=0\quad\text{(in other words $\omega\in{\mathbb{C}}$)}.$ (3.4)
Let us define the differential operator
${\cal C}(\sigma,\kappa)$ $:=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa(z)\big{)}\partial_{z}+\tfrac{1}{2}\kappa^{\prime}$ (3.5)
$=\partial_{z}\sigma(z)\partial_{z}+\tfrac{1}{2}\big{(}\partial_{z}\kappa(z)+\kappa(z)\partial_{z}\big{)}.$
Clearly, the class of operators (3.3) coincides with the class of operators
${\cal C}(\sigma,\kappa)+\omega$ (3.6)
Note that ${\mathcal{C}}(\sigma,\kappa)$ coincides with ${\mathcal{C}}^{0}$, the Casimir operator for Miller’s Lie algebra restricted to the subspace $N=0$, see (2.41).
3.3 Basic symmetry
Recall from (2.43) that $\rho(z)$ is defined as a solution of
$(\sigma(z)\partial_{z}-\kappa(z))\rho(z)=0.$ (3.7)
We have the identity
${\cal C}(\sigma,\kappa)=\rho^{-1}(z)\partial_{z}\sigma(z)\rho(z)\partial_{z}+\tfrac{1}{2}\kappa^{\prime}.$ (3.8)
The following theorem describes a certain symmetry of the entire family of hypergeometric class equations. We will call it the basic symmetry.
Theorem 3.1
We have
$\begin{array}[]{l}\rho(z){\cal C}(\sigma,\kappa)\rho^{-1}(z)={\cal C}(\sigma,-\kappa).\end{array}$ (3.9)
Hence,
${\cal C}(\sigma,\kappa)F=0\ \Rightarrow\ {\cal C}(\sigma,-\kappa)\rho F=0.$ (3.10)
Proof.  We multiply the equation
$0=\left(\rho^{-1}(z)\partial_{z}\sigma(z)\rho(z)\partial_{z}+\tfrac{1}{2}\kappa^{\prime}\right)f(z)$
with $\rho(z)$. We get
$\begin{array}[]{rl}0&=\left(\partial_{z}\sigma(z)\rho(z)\partial_{z}\rho^{-1}(z)+\frac{1}{2}\kappa^{\prime}\right)\rho(z)f(z)\\[8.53581pt] &=(\sigma(z)\partial_{z}^{2}+(-\kappa(z)+\sigma^{\prime}(z))\partial_{z}-\frac{1}{2}\kappa^{\prime})\rho(z)f(z).\end{array}$
$\Box$
Applying twice the basic symmetry we get the identity. Hence we obtain a group of symmetries of hypergeometric class equations isomorphic to ${\mathbb{Z}}_{2}$.
Note that the basic symmetry of Theorem 3.10 corresponds to the symmetry $\tau$ of Miller’s Lie algebra, see (2.21) and Theorem 2.1.
3.4 Basic pair of factorizations
It is often useful to represent a 2nd order operator as a product of two 1st order operators plus a constant. In this section we describe a pair of such factorizations of hypergeometric class operators, which can be easily formulated in a unified way. It is convenient to formulate them for a family indexed by $n\in{\mathbb{C}}$. These factorizations lead to recurrence relations for hypergeometric class functions.
Fix $\sigma$, a polynomial of degree $\leq 2$, $\kappa_{0}$, a polynomial of degree $\leq 1$ and $\omega_{0}\in{\mathbb{C}}$. We set
$\kappa_{n}(z):=$ $n\sigma^{\prime}(z)+\kappa_{0}(z),$ (3.11)
$\omega_{n}:=$ $n^{2}\frac{\sigma^{\prime\prime}}{2}+n\kappa_{0}^{\prime}+\omega_{0}.$ (3.12)
Theorem 3.2
(1) Factorization properties
$\begin{array}[]{rl}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}&=\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\partial_{z}+n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}\\[8.53581pt] &=\partial_{z}\Big{(}\sigma(z)\partial_{z}+\kappa_{n}(z)\Big{)}+n(n-1)\frac{\sigma^{\prime\prime}}{2}+(n-\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}.\end{array}$
(2) Transmutation properties
$\begin{array}[]{rl}\partial_{z}\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}&=\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\big{)}\partial_{z},\\[8.53581pt] \Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}&=\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}.\end{array}$
(3) Recurrence relations
$\begin{array}[]{rll}&\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}F=0&\Rightarrow\qquad\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}\partial_{z}F=0,\\[8.53581pt] &\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}F=0&\Rightarrow\qquad\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}F=0.\end{array}$
The argument used in the proof of the implication (1$\Rightarrow$(2) is typical for the so-called supersymmetric quantum mechanics and is described in the following lemma:
Lemma 3.3
Suppose that for $n\in{\mathbb{C}}$, ${\cal A}_{+}^{n+\frac{1}{2}}$, ${\cal A}_{-}^{n-\frac{1}{2}}$, ${\cal C}^{n}$ are operators and $\eta_{n\pm\frac{1}{2}}$ are numbers satisfying
$\begin{array}[]{rl}{\cal C}^{n}&={\cal A}_{+}^{n+\frac{1}{2}}{\cal A}_{-}^{n-\frac{1}{2}}+\eta_{n-\frac{1}{2}},\\[8.53581pt] &={\cal A}_{-}^{n-\frac{1}{2}}{\cal A}_{+}^{n+\frac{1}{2}}+\eta_{n+\frac{1}{2}}.\end{array}$
Then
$\begin{array}[]{l}{\cal A}_{-}^{n-\frac{1}{2}}{\cal C}^{n}={\cal C}^{n-1}{\cal A}_{-}^{n-\frac{1}{2}},\\[8.53581pt] {\cal A}_{+}^{n+\frac{1}{2}}{\cal C}^{n}={\cal C}^{n+1}{\cal A}_{+}^{n+\frac{1}{2}}\end{array}$
Proof of Theorem 3.2 (1) follows by a direct computation. (1) implies (2) by Lemma 3.3 if we set
$\begin{array}[]{l}{\cal A}_{+}^{n+\frac{1}{2}}=\partial_{z},\ \ \ \ {\cal A}_{-}^{n-\frac{1}{2}}=\sigma(z)\partial_{z}+\kappa_{n}(z),\ \ \ \ \eta_{n+\frac{1}{2}}=n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa^{\prime}+\omega,\\[8.53581pt] {\cal C}^{n}:={\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}.\end{array}$ (3.13)
(2) easily implies (3). $\Box$
Note that ${\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}$ from this section essentially coincides with ${\mathcal{C}}+\omega$ restricted to ${\cal Z}^{n}$ from Section 2.
3.5 Ladders of solutions
Let $\kappa_{n},\omega_{n}$ be as in (3.11) and (3.12). Suppose we have a solution
$\Big{(}{\mathcal{C}}(\sigma,\kappa_{0})+\omega_{0}\Big{)}f=0.$ (3.14)
Then from $f$ we can construct a two-sided ladder of solutions. More precisely, for any $n\in{\mathbb{N}}_{0}$
${\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}\quad\text{annihilates}\quad$ $\partial_{z}^{n}f,$ (3.15)
${\mathcal{C}}(\sigma,\kappa_{-n})+\omega_{-n}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}f.$ (3.16)
Note that (3.15) is related to (3.16) by the basic symmetry (3.10).
To see (3.16), we note that
$\sigma(z)\partial_{z}+\kappa_{-j}(z)$ $=\sigma(z)^{j+1}\rho(z)^{-1}\partial_{z}^{j}\sigma(z)^{-j}\rho(z),$ (3.17)
hence ($\text{hence }\big{(}\sigma(z)\partial_{z}+\kappa_{-n}(z)\big{)}\cdots\big{(}\sigma(z)\partial_{z}+\kappa(z)\big{)}$ $=\sigma(z)^{n}\rho_{0}(z)^{-1}\partial_{z}^{n}\rho_{0}(z).$ (3.18)
Consider now special cases $\omega_{0}=\mp\frac{\kappa_{0}^{\prime}}{2}$:
${\mathcal{C}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z},$ (3.19)
${\mathcal{C}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z}+\kappa_{0}^{\prime}.$ (3.20)
They have elementary solutions
${\mathcal{C}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $1,$ (3.21)
${\mathcal{C}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $\rho_{0}^{-1}.$ (3.22)
(3.21) is obvious. To see (3.22) we differentiate
$(\sigma(z)\partial_{z}+\kappa_{0}(z))\rho_{0}^{-1}(z)=0,$
obtaining
$0=(\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)+\kappa_{0}(z))\partial_{z}+\kappa_{0}^{\prime})\rho_{0}^{-1}(z)\$ (3.23)
Alternatively, we can derive (3.22) from (3.21) by the symmetry (3.10). The special solutions (3.21) and (3.22) lead to a pair of one-sided ladders linked by the symmetry (3.10): for any $n\in{\mathbb{N}}_{0}$
${\mathcal{C}}(\sigma,-n\sigma^{\prime}+\kappa_{0})-(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}$ (3.24)
${\mathcal{C}}(\sigma,n\sigma^{\prime}+\kappa_{0})+(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\partial_{z}^{n}\rho_{0}^{-1}.$ (3.25)
(3.24) consists of polynomials, which we will analyze in more detail in Section 7.
3.6 Chebyshev ladder
Let us fix $\sigma$, as usual a polynomial of degree $\leq 2$, and $\omega\in{\mathbb{C}}$. The equations given by the following two operators can be easily solved in elementary functions:
${\mathcal{C}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}\partial_{z}+\omega$ (3.26)
$=\sigma(z)\partial_{z}^{2}+\frac{\sigma^{\prime}(z)}{2}\partial_{z}+\omega,$ (3.27)
${\mathcal{C}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\partial_{z}\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}+\omega$ (3.28)
$=\sigma(z)\partial_{z}^{2}+\frac{3\sigma^{\prime}(z)}{2}\partial_{z}+\frac{\sigma^{\prime\prime}}{2}+\omega.$ (3.29)
In fact, set
$y(z)=\int_{0}^{z}\frac{{\rm d}x}{\sqrt{\sigma(x)}},$ (3.30)
which solves the equation $\frac{{\rm d}y}{{\rm d}z}=\frac{1}{\sqrt{\sigma(z)}}$. Then
${\mathcal{C}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)},$ (3.31)
${\mathcal{C}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\frac{1}{\sqrt{\sigma(z)}}\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)}.$ (3.32)
We can embed the parameters of (3.26) and (3.28) into a single ladder as follows. We set $\kappa_{0}=0$ and $\omega_{0}=\omega-\frac{\sigma^{\prime\prime}}{8}$, so that
$\kappa_{m}=m\sigma^{\prime},\quad\omega_{m}=\omega+\Big{(}m^{2}-\frac{1}{4}\Big{)}\frac{\sigma^{\prime\prime}}{2}.$ (3.33)
then
[…]
By using the recurrence relations (3.15) and (3.16) we obtain elementary solutions for ${\mathcal{C}}(\sigma,\kappa_{m})+\omega_{m}$ with $m\in{\mathbb{Z}}+\frac{1}{2}$.
4 Singular point
Throughout this section we assume that
$\sigma(0)=0.$ (4.1)
We will discuss solutions of hypergeometric class equations given by power series around 00. We will also describe some symmetries, which exist if (4.1) holds.
4.1 Solutions around a singular point
(4.1) usually means that 00 is a singular point of the equation (3.6). In fact, rewriting the equation (3.6) as
$\frac{1}{\sigma(z)}\big{(}{\mathcal{C}}+\omega\big{)}f(z)=\Big{(}\partial_{z}^{2}+\frac{\tau(z)}{\sigma(z)}\partial_{z}+\frac{\omega}{\sigma(z)}\Big{)}f(x)=0$ (4.2)
we see that $\frac{\tau(z)}{\sigma(z)}$ or $\frac{\omega}{\sigma(z)}$ will usually have a singularity at 00. Straightforward calculations (known under the name of the Frobenius method [WW]) lead then to the following result
Theorem 4.1
There exists a unique formal power series $F(z)$ solving
$\Big{(}{\cal C}(\sigma,\kappa)+\omega\Big{)}F(z)=0,\ \ \ F(0)=1.$ (4.3)
It is equal to
$F(\sigma,\kappa,\omega;z)=\sum_{n=0}^{\infty}\frac{\Pi_{j=0}^{n-1}(\omega+(j+\frac{1}{2})\kappa^{\prime}+j(j+1)\frac{\sigma^{\prime\prime}}{2})}{\Pi_{j=0}^{n-1}(\kappa(0)+(j+1)\sigma^{\prime}(0))n!}(-z)^{n}$ (4.4)
If $\sigma^{\prime}(0)=0$ and $\sigma^{\prime\prime}\neq 0$, then $\sigma(z)$ has a double root at zero and then the series (4.4) is divergent. However, it can be shown to be asymptotic to one of the solutions of (3.6) (see Section 6.3). In all other cases the series (4.4) is convergent. The function $F(\sigma,\kappa,\omega;z)$ given by (4.4) will be called the unified hypergeometric function. It depends meromorphically on $\sigma^{\prime}(0),\sigma^{\prime\prime},\kappa,\omega$.
Recall that $\kappa_{n}$ and $\omega_{n}$ were defined in (3.11) and (3.12). The basic pair of recurrence relations is as follows:
$\partial_{z}F(\sigma,\kappa_{n},\omega_{n};z)$ $=-\frac{\omega_{n}+\frac{1}{2}\kappa_{n}^{\prime}}{\kappa_{n+1}(0)}F(\sigma,\kappa_{n+1},\omega_{n+1};z),$ (4.5)
$\left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)F(\sigma,\kappa_{n+1},\omega_{n+1};z)$ $=\kappa_{n+1}(0)F(\sigma,\kappa_{n},\omega_{n};z).$ (4.6)
4.2 Olver’s normalization
Assume now $\sigma(0)\neq 0$. We set $m:=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. Without sacrificing the generality we can suppose additionally that $\sigma^{\prime}(0)=1$. Thus
$\sigma(z)=z+\frac{\sigma^{\prime\prime}}{2}z^{2},\qquad\kappa(z)=m+\kappa^{\prime}z.$ (4.7)
00 is a regular singular point and 00, $-m$ are its indices [WW].
In this case often instead of the function (4.4) it is more convenient to use the function
${\mathbf{F}}(\sigma,\kappa,\omega;z):=\frac{F(\sigma,\kappa,\omega;z)}{\Gamma(1+m)}=\sum_{n=0}^{\infty}\frac{\Pi_{j=0}^{n-1}(\omega+(j+\frac{1}{2})\kappa^{\prime}+j(j+1)\frac{\sigma^{\prime\prime}}{2})}{\Gamma(m+n+1)n!}(-z)^{n}.$ (4.8)
Note that ${\mathbf{F}}$ is holomorphic in $\sigma^{\prime\prime},\kappa,\omega$. Dividing by $\Gamma(1+m)$, as in (4.8), is sometimes called Olver’s normalization. Here is the basic pair of recurrence relations for the Olver normalized unified hypergeometric function:
$\partial_{z}{\mathbf{F}}(\sigma,\kappa_{n},\omega_{n};z)$ $=-\Big{(}\omega_{n}+\frac{1}{2}\kappa_{n}^{\prime}\Big{)}{\mathbf{F}}(\sigma,\kappa_{n+1},\omega_{n+1};z),$ (4.9)
$\left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right){\mathbf{F}}(\sigma,\kappa_{n+1},\omega_{n+1};z)$ $={\mathbf{F}}(\sigma,\kappa_{n},\omega_{n};z).$ (4.10)
4.3 Power symmetry
We keep the assumptions of Subsection 4.2. Let us introduce a certain transformation of the parameters. We keep $\sigma$ the same, and the transformed $\kappa$ and $\omega$ are
$\kappa^{\circ}(z):=-m+(\kappa^{\prime}-m\sigma^{\prime\prime})z,\quad\omega^{\circ}:=\omega-m\kappa^{\prime}+m^{2}\frac{\sigma^{\prime\prime}}{2}.$ (4.11)
Consequently, $m^{\circ}=-m$.
Theorem 4.2
$z^{m}\Big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\Big{)}z^{-m}={\mathcal{C}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.12)
Proof.  Using
$z^{m}\partial_{z}z^{-m}=\partial_{z}-\frac{m}{z},\qquad z^{m}\partial_{z}^{2}z^{-m}=\partial_{z}^{2}-\frac{2m}{z}\partial_{z}+\frac{m(m+1)}{z^{2}}$ (4.13)
we compute
$z^{m}\Big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\Big{)}z^{-m}$ (4.14)
$=$ $z^{m}\Big{(}\Big{(}z+\frac{\sigma^{\prime\prime}}{2}z^{2}\Big{)}\partial_{z}^{2}+\big{(}1+m+(\sigma^{\prime\prime}+\kappa^{\prime})z\big{)}\partial_{z}+\frac{\kappa^{\prime}}{2}+\omega\Big{)}z^{-m}$ (4.15)
$=$ $\Big{(}z+\frac{\sigma^{\prime\prime}}{2}z^{2}\Big{)}\partial_{z}^{2}+\Big{(}1-m+\big{(}(1-m)\sigma^{\prime\prime}+\kappa^{\prime}\big{)}z\Big{)}\partial_{z}+\frac{\sigma^{\prime\prime}}{2}m(m-1)-\kappa^{\prime}(m-\tfrac{1}{2})+\omega$ (4.16)
$=$ ${\mathcal{C}}(\sigma,\kappa^{\circ})+\omega^{\circ}.$ (4.17)
Note that $\kappa^{\circ\circ}=\kappa$ and $\omega^{\circ\circ}=\omega$. Thus we obtain a ${\mathbb{Z}}_{2}$ symmetry of hypergeometric class equations, different from the basic symmetry of Thm 3.10.
As a consequence of (4.12), both $\mathbf{F}(\sigma,\kappa,\omega;z)$ and $z^{-m}{\mathbf{F}}(\sigma,\kappa^{\circ},\omega^{\circ};z)$ are annihilated by ${\mathcal{C}}(\sigma,\kappa)+\omega$. If $m\not\in{\mathbb{Z}}$ they form a basis of solutions of the corresponding equation. The situation for $m\in{\mathbb{Z}}$ will be discussed in Section 8 about the degenerate case.
4.4 Inversion symmetry
We still assume $\sigma(0)=0$. Besides, we suppose that
$\sigma^{\prime\prime}\neq 0\quad\text{ or }\quad\kappa^{\prime}\neq 0.$ (4.18)
Suppose that $\zeta\in{\mathbb{C}}$ solves the equation
$\frac{\sigma^{\prime\prime}}{2}\zeta^{2}+(\sigma^{\prime\prime}+\kappa^{\prime})\zeta+\frac{\kappa}{2}+\omega=0.$ (4.19)
Note that (4.18) guarantees that (4.19) has a solution. Set
$\sigma^{\triangle}$ $:=\frac{\sigma^{\prime\prime}}{2}w-\sigma^{\prime}(0)w^{2},$ (4.20)
$\kappa^{\triangle}$ $=\big{(}\kappa(0)+2(\zeta+1)\sigma^{\prime}(0)\big{)}w-\sigma^{\prime\prime}(1+\zeta)-\kappa^{\prime},$ (4.21)
$\omega^{\triangle}$ $:=-\sigma^{\prime}(0)(1+\zeta)^{2}-\kappa(0)(\zeta+\tfrac{1}{2}).$ (4.22)
Theorem 4.3
Consider the substitution $w=-\frac{1}{z}$. Then we have
$-z^{-\zeta+1}\Big{(}{\mathcal{C}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}z^{\zeta}={\mathcal{C}}(\sigma^{\triangle},\kappa^{\triangle};w,\partial_{w})+\omega^{\triangle}$ (4.23)
Hence
${\mathcal{C}}(\sigma,\kappa)+\omega\quad\text{ annihilates }\quad z^{-\zeta}\mathbf{F}(\sigma^{\triangle},\kappa^{\triangle},\omega^{\triangle};-z^{-1}).$ (4.24)
Proof.  Indeed, using $\partial_{z}=w^{2}\partial_{w}$ and $\partial_{z}^{2}=w^{4}\partial_{w}^{2}+2w^{3}\partial_{w}$, we obtain
$-z\Big{(}{\mathcal{C}}(\sigma,\kappa;z,\partial_{z})+\omega\Big{)}$ (4.25)
$=$ $\sigma(-w^{-1})w^{3}\partial_{w}^{2}+\Big{(}2\sigma(-w^{-1})w^{2}+\sigma^{\prime}(-w^{-1})w+\kappa(-w^{-1})w\Big{)}\partial_{w}+\Big{(}\frac{\kappa^{\prime}}{2}+\omega\Big{)}w^{-1}.$ (4.26)
Then we use $w^{\zeta}\partial_{w}w^{-\zeta}=\partial_{w}-\frac{\zeta}{w}$ and $w^{\zeta}\partial_{w}^{2}w^{-\zeta}=\partial_{w}^{2}-\frac{2\zeta}{w}\partial_{w}+\frac{(\zeta+\zeta^{2})}{w^{2}}$ to obtain
[…]
$=$ $\Big{(}\frac{\sigma^{\prime\prime}}{2}w-\sigma^{\prime}(0)w^{2}\Big{)}\partial_{w}^{2}+\Big{(}\big{(}\kappa(0)+2\zeta\sigma^{\prime}(0)\big{)}w-\frac{\sigma^{\prime\prime}}{2}-\zeta\sigma^{\prime\prime}-\kappa^{\prime}\Big{)}\partial_{w}$ (4.28)
$-\sigma^{\prime}(0)(\zeta+\zeta^{2})-\kappa(0)\zeta+\Big{(}\frac{\kappa^{\prime}}{2}+\omega+\frac{\sigma^{\prime\prime}}{2}(2\zeta+\zeta^{2})+\kappa^{\prime}\zeta\Big{)}w^{-1}$ (4.29)
$=$ $\sigma^{\triangle}(w)\partial_{w}^{2}+\big{(}{\sigma^{\triangle}}^{\prime}(w)+\kappa^{\triangle}(w)\big{)}\partial_{w}+\frac{{\kappa^{\triangle}}^{\prime}}{2}+\omega^{\triangle}.$ (4.30)
$\Box$
Note that $\sigma^{\triangle\triangle}=\sigma$, $\kappa^{\triangle\triangle}=\kappa$, $\omega^{\triangle\triangle}=\omega$. Hence the inversion symmetry generates a group isomorphic to ${\mathbb{Z}}_{2}$.
Remark 4.4
It is easy to see that the substitution $z\mapsto\frac{1}{w}$ also generates a symmetry of hypergeometric class equations. This follows immediately from Theorem 4.24 and the trivial fact that $z\mapsto-z$ leads to a symmetry as well.
5 Integral representations
Hypergeometric class functions possess useful integral representations. Most of them have the form of an Euler transform of an elementary function. In Subsection 5.1 we will show how to derive these representaions in a unified way. Note that typically for a given equation one can choose various contours of integration, obtaining various solutions.
Some important integral represenations have a different form—they can be viewed as Laplace transforms of certain elementary functions. They will be described in Subsection 5.2.
In this section we will often deal with multivalued functions $s\mapsto f(s)$ defined on a certain Riemann surface. These functions are analytically continued along a certain curve $\gamma$ contained in this Riemann surface. We will use the notation
$f(s)\Big{|}_{s_{0}}^{s_{1}}=f(s_{1})-f(s_{0}),$ (5.1)
where $s_{1}$ and $s_{2}$ are the endpoints of the curve $\gamma$.
As usual, we fix $\sigma,\kappa_{0}$, and for $n\in{\mathbb{C}}$, as in (3.11), we set
$\kappa_{n}(z):=n\sigma^{\prime}(z)+\kappa_{0}(z).$
In this section $\omega_{0}=\frac{\kappa_{0}^{\prime}}{2}$, so that according to (3.12),
$\omega_{n}:=n^{2}\frac{\sigma^{\prime\prime}}{2}+(n+\tfrac{1}{2})\kappa_{0}^{\prime}.$ (5.2)
5.1 Euler transforms
As usual, we assume that
$(\sigma(z)\partial_{z}-\kappa_{0}(z))\rho_{0}(z)=0.$ (5.3)
Theorem 5.1
[…]
$\sigma(s)(s-z)^{-n-2}\rho_{0}^{-1}(s)\Big{|}_{{\rm s}(0)}^{s(1)}=0$ (5.4)
and
$f_{n}(z):=\int_{\gamma}(s-z)^{-n-1}\rho_{0}^{-1}(s){\rm d}s$ (5.5)
is well-defined. Then $f_{n}$ is annihillated by
${\cal C}(\sigma,\kappa_{n})+\omega_{n}$ (5.6)
$=$ $\sigma(z)\partial_{z}^{2}+(\kappa_{0}(z)+(n+1)\sigma^{\prime}(z))\partial_{z}+\kappa_{0}^{\prime}(n+1)+\frac{\sigma^{\prime\prime}}{2}n(n+1).$ (5.7)
Besides, if $f_{n+1}$ is also well-defined, then
$\partial_{z}f_{n}(z)$ $=(n+1)f_{n+1}(z)$ (5.8)
$\left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)f_{n+1}(z)$ $=-\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)f_{n}(z)$ (5.9)
Proof.  For simplicity, we will assume that both $f_{n}$ and $f_{n+1}$ are well defined. (5.8) is obvious. Let us prove (5.9). Using the Taylor expansion of the polynomials $\sigma(z)$ and $\kappa(z)$ we obtain
$\Big{(}\sigma(z)\partial_{z}+\kappa_{0}(z)+(n+1)\sigma^{\prime}(z)\Big{)}f_{n+1}(z)$ (5.10)
$=$ $(n+2)\int_{\gamma}\Big{(}\sigma(s)+\sigma^{\prime}(s)(z-s)+\frac{\sigma^{\prime\prime}}{2}(z-s)^{2}\Big{)}(s-z)^{-n-3}\rho_{0}(s)^{-1}{\rm d}s$ (5.11)
$+\int_{\gamma}\Big{(}\kappa_{0}(s)+(z-s)\kappa^{\prime}+(n+1)\sigma^{\prime}(s)+(n+1)(z-s)\sigma^{\prime\prime}\Big{)}(s-z)^{-n-2}\rho_{0}(s)^{-1}{\rm d}s$ (5.12)
$=$ $-\Big{(}n\frac{\sigma^{\prime\prime}}{2}+\kappa_{0}^{\prime}\Big{)}\int_{\gamma}(s-z)^{-n-1}\rho_{0}(s)^{-1}{\rm d}s$ (5.13)
$+\int_{\gamma}\Big{(}(n+2)\sigma(s)(s-z)^{-n-3}-\sigma^{\prime}(s)(s-z)^{-n-2}+\kappa_{0}(s)(s-z)^{-n-2}\Big{)}\rho_{0}(s)^{-1}{\rm d}s$ (5.14)
$=$ $-\Big{(}n\frac{\sigma^{\prime\prime}}{2}+\kappa_{0}^{\prime}\Big{)}f_{n}(z)$ (5.15)
$+\int_{\gamma}\Big{(}\partial_{s}\sigma(s)(s-z)^{-n-2}\rho_{0}(s)^{-1}\Big{)}{\rm d}s$ (5.16)
$+\int_{\gamma}(s-z)^{-n-2}\Big{(}\sigma(s)\partial\rho_{0}(s)^{-1}+\kappa_{0}(s)\rho_{0}(s)^{-1}\Big{)}{\rm d}s.$ (5.17)
(5.16) vanishes because of (5.4). (5.17) is zero by (5.3). $\Box$
It is sometimes useful to consider differently normalized Euler integrals:
$\mathbf{f}_{n}(z):=\frac{f_{n}(z)}{\Gamma(n+1)}=\frac{1}{\Gamma(n+1)}\int_{\gamma}(s-z)^{-n-1}\rho_{0}^{-1}(s){\rm d}s$ (5.18)
Then the recurrence relations (5.8) and (5.9) are modified:
$\begin{array}[]{rl}\partial_{z}\mathbf{f}_{n}(z)&=\mathbf{f}_{n+1}(z)\\[5.69054pt] \left(\sigma(z)\partial_{z}+\kappa_{n+1}(z)\right)\mathbf{f}_{n+1}(z)&=-(n+1)\left(\kappa_{0}^{\prime}+\frac{\sigma^{\prime\prime}}{2}n\right)\mathbf{f}_{n}(z)\end{array}$ (5.19)
5.2 Laplace integrals
Suppose that $\sigma^{\prime\prime}=0$. We still assume (5.2), which now can be rewritten as
$\omega_{n}=(n+\tfrac{1}{2})\kappa_{0}^{\prime}.$ (5.20)
Let
$\Big{(}\sigma(-\partial_{s})s+\kappa_{0}(-\partial_{s})\Big{)}\delta_{0}(s)=0,$ (5.21)
which defines up to a coefficient an elementary function $\delta_{0}$.
Theorem 5.2
[…]
$(s^{n+2}\sigma^{\prime}+s^{n+1}\kappa^{\prime})\delta_{0}(s){\rm e}^{sz}\Big{|}_{s(0)}^{s(1)}=0,$ (5.22)
and the following integral exists:
$g_{n}(z)=\int_{\gamma}\delta_{0}(s)s^{n}{\rm e}^{zs}{\rm d}s.$ (5.23)
Then $g_{n}$ is annihillated by
${\cal C}(\sigma,\kappa_{n})+\omega_{n}$ (5.24)
=\,= $\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)(n+1)+\kappa_{0}(z))\partial_{z}+(n+1)\kappa_{0}^{\prime}.$ (5.25)
Besides, if also $g_{n+1}$ is well defined, then
$\partial_{z}g_{n}(z)$ $=g_{n+1},$ (5.26)
$(\sigma(z)\partial_{z}+\kappa_{n+1}(z))g_{n+1}(z)$ $=-(n+1)\kappa_{0}^{\prime}g_{n}(z).$ (5.27)
Proof.  For simplicity, we will assume that both $g_{n}$ and $g_{n+1}$ are well defined, and we will prove the recurrence relations, which directly imply (5.25).
(5.26) is obvious. Let us show (5.27):
$(\sigma(z)\partial_{z}+(n+1)\sigma^{\prime}+\kappa_{0}(z))g_{n+1}$ (5.28)
$=$ $\int_{\gamma}\Big{(}\sigma^{\prime}\delta_{0}(s)s^{n+2}z+\sigma(0)\delta_{0}(s)s^{n+2}+(n+1)\sigma^{\prime}\delta_{0}(s)s^{n+1}\Big{)}{\rm e}^{zs}{\rm d}s$ (5.29)
$+\int_{\gamma}\Big{(}\kappa_{0}^{\prime}\delta_{0}(s)s^{n+1}z+\kappa_{0}(0)\delta_{0}(s)s^{n+1}\Big{)}{\rm e}^{zs}{\rm d}s$ (5.30)
$=$ $\int_{\gamma}{\rm d}s\partial_{s}\Big{(}\sigma^{\prime}\delta_{0}(s)s^{n+2}{\rm e}^{zs}+\kappa_{0}^{\prime}\delta_{0}(s)s^{n+2}{\rm e}^{zs}\Big{)}$ (5.31)
$+\int_{\gamma}\Big{(}-\sigma^{\prime}s\delta_{0}^{\prime}(s)-\sigma^{\prime}\delta_{0}(s)-\kappa_{0}^{\prime}\delta_{0}^{\prime}(s)+\sigma(0)s\delta_{0}(s)+\kappa_{0}(0)\delta_{0}(s)\Big{)}s^{n+1}{\rm e}^{zs}{\rm d}s$ (5.32)
$-(n+1)\kappa_{0}^{\prime}\int\delta_{0}(s)s^{n}{\rm e}^{zs}{\rm d}s.$ (5.33)
Now (5.31) vanishes because of (5.22) and (5.32) due to (5.21). $\Box$
Sometimes it will be more convenient to present the integral represention (5.23) in a different form, with the variable $s$ replaced with $t=\frac{1}{s}$:
$g_{n}(z)=\int_{\tilde{\gamma}}\delta_{0}(t^{-1})t^{-n+2}{\rm e}^{\frac{z}{t}}{\rm d}t.$ (5.34)
where the contour $\tilde{\gamma}$ is obtaned by the change of variable and reversing the orientation.
6 Applications case by case
In this section we consider all five nontrivial types of hypergeometric equations, listed in the introduction. We describe those of their properties that are direct consequences of the unified theory, discussed in the previous sections. We try to follow the same pattern. Here is the list of items that we will give for each hypergeometric type, if available:
  • Operator that generates the equation, that is ${\mathcal{C}}(\sigma,\kappa)+\omega$, expressed in the traditional form.
  • Parameters $\sigma(z),\kappa(z),\omega$; see (3.4) and (3.5).
  • Weight $\rho(z)$, see (2.43) or (3.7).
  • Basic symmetry, see (3.9).
  • Power symmetry (if available), see (4.12).
  • Inversion symmetry (if available), see (4.24).
  • Theorem about integral representations, see Theorems 5.1 and 5.2.
  • Standard solution annihillated by ${\mathcal{C}}(\sigma,\kappa)+\omega$ (typically a special case of the unified hypergeometric function (4.4), but for the Hermite equation we need to make an exception).
  • Olver normalized standard solution (if available), see (4.8).
  • Basic pair of recurrence relations for standard solutions, see (4.5)-(4.6) or (4.9)-(4.10).
  • Integral representations of standard solution (obtained by an application of Theorems 5.1 or 5.2).
  • Chebyshev solutions (if available), see Subsection 3.6.
We will often use the Pochhammer symbols
$(a)_{j}:=a(a+1)\cdots(a+j-1),\qquad a\in{\mathbb{C}},\quad j\in{\mathbb{N}}_{0}.$ (6.1)
6.1 The ${}_{2}F_{1}$ equation
${}_{2}F_{1}$ operator:
$\mathcal{F}(a,b;c):=z(1-z)\partial_{z}^{2}+\big{(}c-(a+b+1)z\big{)}\partial_{z}-ab.$ (6.2)
Parameters:
$\sigma(z)=z(1-z),\quad\kappa(z)=c-1-(a+b-1)z,\quad\omega=-\big{(}a-\tfrac{1}{2}\big{)}\big{(}b-\tfrac{1}{2}\big{)}-\tfrac{1}{4}.$
Weight:
$\rho(z)=z^{c-1}(z-1)^{a+b-c}.$
Basic symmetry:
$z^{c-1}(z-1)^{a+b-c}{\cal F}(a,b;c)z^{1-c}(z-1)^{-a-b+c}={\cal F}(1-b,1-a;2-c).$
Power symmetry:
$z^{c-1}{\cal F}(a,b;c)z^{1-c}={\cal F}(b+1-c,a+1-c;2-c).$
Inversion symmetry:
$(-z)^{1+a}{\cal F}(a,b;c;z,\partial_{z})(-z)^{-a}={\cal F}(a,a-c+1;a-b+1;w,\partial_{w}),\qquad w=z^{-1}.$
Theorem 6.1 (about integral representations)
[…]
$t^{b-c+1}(1-t)^{c-a}(t-z)^{-b-1}\Big{|}_{\gamma(0)}^{\gamma(1)}=0.$
Then
${\cal F}(a,b;c)\quad\text{ annihilates }\quad\int_{\gamma}t^{b-c}(1-t)^{c-a-1}(t-z)^{-b}{\rm d}t.$ (6.3)
Proof.  We check that
${\cal F}(a,b;c)t^{b-c}(1-t)^{c-a-1}(t-z)^{-b}=-b\partial_{t}t^{b-c+1}(1-t)^{c-a}(t-z)^{-b-1}.$ (6.4)
$\Box$
${}_{2}F_{1}$ function: For $|z|<1$ defined by the following power series, then analytically extended:
$F(a,b;c;z)=\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{(c)_{j}}\frac{z^{j}}{j!},\quad|z|<1.$
Olver normalized ${}_{2}F_{1}$ function:
${\bf F}(a,b;c;z):=\frac{F(a,b,c,z)}{\Gamma(c)}=\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{\Gamma(c+j)}\frac{z^{j}}{j!}.$
Integral representation:
$\int_{1}^{\infty}t^{b-c}(t-1)^{c-a-1}(t-z)^{-b}{\rm d}t$
$=$ $\Gamma(a)\Gamma(c-a)\mathbf{F}(a,b;c;z),\ \ \ \ \mathrm{Re}(c-a)>0,\ \mathrm{Re}a>0,\ \ \ z\not\in[1,\infty[.$
Basic pair of recurrence relations:
$\partial_{z}\mathbf{F}(a,b;c;z)$ $=ab\mathbf{F}(a+1,b+1;c+1;z),$ (6.6)
$\Big{(}z(1-z)\partial_{z}+\big{(}c-(a+b+1)z\big{)}\Big{)}\mathbf{F}(a+1,b+1;c+1;z)$ $=\mathbf{F}(a,b;c;z).$ (6.7)
Chebyshev solutions for $k=0,1,\dots$:
${\bf F}\Big{(}1+k+\lambda,1+k-\lambda;\frac{3}{2}+k;\frac{1-w}{2}\Big{)}$ $=\frac{(1-w^{2})^{\frac{1}{2}+k}}{2\sqrt{\pi}(-2)^{k}}\partial_{w}^{k}\frac{\big{(}w+\mathrm{i}\sqrt{1-w^{2}}\big{)}^{\lambda}+\big{(}w-\mathrm{i}\sqrt{1-w^{2}}\big{)}^{\lambda}}{\sqrt{1-w^{2}}},$ (6.8)
${\bf F}\Big{(}-k+\lambda,-k-\lambda;\frac{1}{2}-k;\frac{1-w}{2}\Big{)}=$ $\frac{2^{k}}{\mathrm{i}\sqrt{\pi}(\lambda-k)_{2k+1}}\partial_{w}^{k}\frac{\big{(}w+\mathrm{i}\sqrt{1-w^{2}}\big{)}^{\lambda}-\big{(}w-\mathrm{i}\sqrt{1-w^{2}}\big{)}^{\lambda}}{\sqrt{1-w^{2}}}.$ (6.9)
6.2 The ${}_{1}F_{1}$ equation
${}_{1}F_{1}$ operator
${\cal F}(a;c):=z\partial_{z}^{2}+(c-z)\partial_{z}-a.$ (6.10)
Parameters
$\sigma(z)$ $=z,\quad\kappa(z)=c-1-z,\quad\omega=-a+\tfrac{1}{2}.$
Weight:
$\rho(z)={\rm e}^{-z}z^{c-1}.$
Theorem 6.2 (about integral representations)
  1. 1.
[…]
    $t^{a-c+1}{\rm e}^{t}(t-z)^{-a-1}\Big{|}_{t(0)}^{t(1)}=0.$
    Then
    ${\cal F}(a;c)\quad\text{ annihilates }\quad\int_{\gamma}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t.$ (6.11)
  2. 2.
[…]
    ${\rm e}^{\frac{z}{t}}t^{-c}(1-t)^{c-a}\Big{|}_{t(0)}^{t(1)}=0.$
    Then
    ${\cal F}(a;c)\quad\text{ annihilates }\quad\int_{\gamma}{\rm e}^{\frac{z}{t}}t^{-c}(1-t)^{c-a-1}{\rm d}t.$ (6.12)
Proof.
${\cal F}(a,c)\,t^{a-c}{\rm e}^{t}(t-z)^{-a}$ $=-a\partial_{t}t^{a-c+1}{\rm e}^{t}(t-z)^{-a-1},$ (6.13)
${\cal F}(a,c)\,{\rm e}^{\frac{z}{t}}t^{-c}(1-t)^{c-a-1}$ $=-\partial_{t}{\rm e}^{\frac{z}{t}}t^{-c}(1-t)^{c-a}.$ (6.14)
$\Box$
Basic symmetry:
$z^{c-1}{\rm e}^{-z}{\cal F}(a;c;z,\partial_{z})z^{1-c}{\rm e}^{z}={\cal F}(1-a;2-c;w,\partial_{w}),\quad w=-z.$
Power symmetry:
$z^{c-1}{\cal F}(a;c)z^{1-c}={\cal F}(1+a-c;2-c).$
Inversion symmetry:
$z^{a+1}{\cal F}(a;c;z,\partial_{z})z^{-a}={\cal F}(a,1+a-c;_{;}w,\partial_{w}),\qquad w=-z^{-1}.$
(Thus the ${}_{1}F_{1}$ equation is equivalent to the ${}_{2}F_{0}$ equation.)
${}_{1}F_{1}$ function:
$F(a;c;z):=\sum_{j=0}^{\infty}\frac{(a)_{j}}{(c)_{j}}\frac{z^{j}}{j!}.$
Olver normalized ${}_{1}F_{1}$ function
${\bf F}(a;c;z):=\frac{F(a;c;z)}{\Gamma(c)}=\sum_{j=0}^{\infty}\frac{(a)_{j}}{\Gamma(c+j)}\frac{z^{j}}{j!}.$
Basic pair of recurrence relations:
$\partial_{z}\mathbf{F}(a;c;z)$ $=a\mathbf{F}(a+1;c+1;z),$ (6.15)
$\big{(}\partial_{z}+c-z)\big{)}\mathbf{F}(a+1;c+1;z)$ $=\mathbf{F}(a;c;z).$ (6.16)
Integral representations: for all parameters
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,(0,z)^{+},-\infty[}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t$ $=$ ${\bf F}(a;c;z),$
for $\mathrm{Re}a>0,\ \mathrm{Re}(c-a)>0$
$\int\limits_{[1,+\infty[}{\rm e}^{\frac{z}{t}}t^{-c}(t-1)^{c-a-1}{\rm d}t$ $=$ $\Gamma(a)\Gamma(c-a)\mathbf{F}(a;c;z),$
6.3 The ${}_{2}F_{0}$ equation
The ${}_{2}F_{0}$ operator:
${{\cal F}}(a,b;-):=z^{2}\partial_{z}^{2}+(-1+(1+a+b)z)\partial_{z}+ab,$ (6.17)
Parameters:
$\sigma(z)=z^{2},\quad\kappa(z)=-1+(a+b-1)z,\quad\omega=\big{(}a-\tfrac{1}{2}\big{)}\big{(}b-\tfrac{1}{2}\big{)}+\tfrac{1}{4}.$
Weight:
$\rho(z)=z^{-1+a+b}{\rm e}^{\frac{1}{z}}.$
Basic symmetry:
$z^{-1+a+b}{\rm e}^{\frac{1}{z}}{\cal F}(a,b;-;z,\partial_{z})z^{1-a-b}{\rm e}^{-\frac{1}{z}}={\cal F}(1-b,1-a;-;w,\partial_{w}),\quad w=-z.$
Inversion symmetry:
$z^{a+1}{\cal F}(a,b;-;z,\partial_{z})z^{-a}=-{\cal F}(a;1+a-b;w,\partial_{w}),\quad w=-z^{-1}.$
Theorem 6.3 (about integral representations)
[…]
${\rm e}^{-\frac{1}{t}}t^{b-a-1}(t-z)^{-b-1}\Big{|}_{t(0)}^{t(1)}=0.$
Then
${\cal F}(a,b;-)\quad\text{ annihilates }\quad\int_{\gamma}{\rm e}^{-\frac{1}{t}}t^{b-a-1}(t-z)^{-b}{\rm d}t.$ (6.18)
A second integral representation is obtained if we interchange $a$ and $b$.
Proof.  We check that
${\cal F}(a,b;-){\rm e}^{-\frac{1}{t}}t^{b-a-1}(t-z)^{-b}=-b\partial_{t}{\rm e}^{-\frac{1}{t}}t^{b-a-1}(t-z)^{-b-1}.$ (6.19)
$\Box$
${}_{2}F_{0}$ function: It is defined for $z\in{\mathbb{C}}\backslash[0,+\infty[$,
$F(a,b;-;z):=\lim_{c\to\infty}F(a,b;c;cz),$
where $|\arg c-\pi|<\pi-\epsilon$, $\epsilon>0$. It extends to an analytic function on the universal cover of \$\0{\mathbb{C}}\backslash\{0\}$ with a branch point of an infinite order at 0. It is annihilated by ${\cal F}(a,b;-)$.
Asymptotic expansion:
$F(a,b;-;z)\sim\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{j!}z^{j},\ |\arg z-\pi|<\pi-\epsilon.$
Basic pair of recurrence relations:
$\partial_{z}F(a,b;-;z)$ $=abF(a+1,b+1;-;z)$ (6.20)
$\Big{(}z^{2}\partial_{z}+\big{(}-1+(1+a+b)z)\big{)}\Big{)}F(a+1,b+1;-;z)$ $=F(a,b;-;z).$ (6.21)
Integral representation for $\mathrm{Re}a>0$:
$\int_{0}^{\infty}{\rm e}^{-\frac{1}{t}}t^{b-a-1}(t-z)^{-b}{\rm d}t=\Gamma(a)F(a,b;-;z),\ \ z\not\in[0,\infty[.$
6.4 The ${}_{0}F_{1}$ equation
${}_{0}F_{1}$ operator:
${\cal F}(c;z,\partial_{z}):=z\partial_{z}^{2}+c\partial_{z}-1.$
Parameters:
$\sigma(z)=z,\quad\kappa(z)=c-1,\quad\omega=-1.$
Weight:
$\rho(z)=z^{c-1}.$
Basic and power symmetry:
$z^{c-1}{\cal F}(c)z^{1-c}$ $={\cal F}(2-c).$ (6.22)
Theorem 6.4 (about integral representations)
Suppose that $[0,1]\ni t\mapsto\gamma(t)$ satisfies
${\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}\Big{|}_{\gamma(0)}^{\gamma(1)}=0.$
Then
${\cal F}(c;z,\partial_{z})\quad\text{ annihilates }\quad\int_{\gamma}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}{\rm d}t.$ (6.23)
Proof.  We check that
${\cal F}(c)\,{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}=-\partial_{t}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}.$ (6.24)
$\Box$
${}_{0}F_{1}$ function
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!}$
Olver normalized ${}_{0}F_{1}$ function:
${\bf F}(c;z):=\frac{F(c;z)}{\Gamma(c)}=\sum_{j=0}^{\infty}\frac{1}{\Gamma(c+j)}\frac{z^{j}}{j!}.$
Basic pair of recurrence relations:
$\partial_{z}\mathbf{F}(c;z)$ $=\mathbf{F}(c+1;z)$ (6.25)
$(\partial_{z}+c)\mathbf{F}(c+1;z)$ $=\mathbf{F}(a;c;z).$ (6.26)
Integral representation for all parameters:
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,0^{+},-\infty[}{\rm e}^{t}{\rm e}^{\frac{z}{t}}t^{-c}{\rm d}t$ $=$ ${\bf F}(c;z),\ \ \ \ \mathrm{Re}z>0;$
Chebyshev solutions for $k=0,1,2,\dots$:
${\bf F}(\tfrac{3}{2}+k;z)$ $=\frac{2^{2k}}{\sqrt{\pi}}\partial_{z}^{k}\frac{\sinh 2\sqrt{z}}{\sqrt{z}},$ (6.27)
${\bf F}(\tfrac{1}{2}-k;z)$ $=\frac{z^{\frac{1}{2}+k}}{\sqrt{\pi}}\partial_{z}^{k}\frac{\cosh 2\sqrt{z}}{\sqrt{z}}.$ (6.28)
6.5 Hermite equation
Hermite operator:
${\cal S}({a}):=\partial_{z}^{2}-2z\partial_{z}-2{a}.$ (6.29)
Parameters:
$\sigma(z)$ $=1,\quad\kappa(z)=-2z,\quad\omega=-2a+1.$
Weight:
$\rho(z)={\rm e}^{-z^{2}}.$
Basic symmetry:
${\rm e}^{-z^{2}}{\cal S}(a;z,\partial_{z}){\rm e}^{z^{2}}=-{\cal S}(1-a;w,\partial_{w}),\qquad w=\pm{\rm i}z.$
Theorem 6.5 (about integral representations)
  1. 1.
    Let $[0,1]\ni t\mapsto\gamma(t)$ satisfy
    ${\rm e}^{t^{2}}(t-z)^{-{a}-1}\Big{|}_{\gamma(0)}^{\gamma(1)}=0.$
    Then
    ${\cal S}({a})\quad\text{annihilates}\quad\int_{\gamma}{\rm e}^{t^{2}}(t-z)^{-{a}}{\rm d}t.$ (6.30)
  2. 2.
    Let $[0,1]\ni t\mapsto\gamma(t)$ satisfy
    ${\rm e}^{-t^{2}-2zt}t^{{a}}\Big{|}_{\gamma(0)}^{\gamma(1)}=0.$
    Then
    ${\cal S}({a})\quad\text{annihilates}\quad\int_{\gamma}{\rm e}^{-t^{2}-2zt}t^{a-1}{\rm d}t.$ (6.31)
Proof.  We check that
${\cal S}({a})\,{\rm e}^{t^{2}}(t-z)^{-{a}}=$ $-{a}\partial_{t}{\rm e}^{t^{2}}(t-z)^{-{a}-1}$
${\cal S}({a})\,{\rm e}^{-t^{2}-2zt}t^{a-1}=$ $-2\partial_{t}{\rm e}^{-t^{2}-2zt}t^{{a}}.$
$\Box$
Hermite function:
$S(a;z)$ ::=:= $z^{-{a}}F\Big{(}\frac{{a}}{2},\frac{{a}+1}{2};-;-z^{-2}\Big{)}.$ (6.32)
Basic pair of recurrence relations:
$\partial_{z}S(a;z)$ $=-aS(a+1;z),$ (6.33)
$\big{(}\partial_{z}-2z\big{)}S(a+1;z)$ $=S(a;z).$ (6.34)
Integral representation for $z\not\in]-\infty,0]$. For $-\frac{1}{2}<{\mathrm{Re}}\omega$:
$\int\limits_{0}^{\infty}{\rm e}^{-t^{2}-2tz}t^{a-1}{\rm d}t$ $=$ $2^{-a}\Gamma(a)S(a;z),$
and for all parameters:
$-{\rm i}\int\limits_{]-{\rm i}\infty,z^{-},{\rm i}\infty[}{\rm e}^{t^{2}}(z-t)^{-a}{\rm d}t$ $=$ $\sqrt{\pi}S(a;z).$
6.6 Lie algebras of symmetries type by type
Miller’s Lie algebra is responsible only for one pair of recurrence relations of hypergeometric class functions. It leads also to a group of discrete symmetries ${\mathbb{Z}}_{2}$. Some of types from the hypergeometric class possess larger sets of recurrence relations, and larger groups of discrete symmetries, which were described e.g. in [De1, De2]. They also possess larger Lie algebras of symmetries, described e.g. in [De2].
In the following table we list all nontrivial types of hypergeometric class equations. We include the Gegenbauer equation, which is the subtype of the ${}_{2}F_{1}$ type with the mirror symmetry $z\mapsto-z$. In this paper there has been no need to consider it separately, however it has special symmetry properties.
The second and third column are based on Section 2 of this paper. The fourth and fifth column are based on [De2]. They describe a more complete Lie algebra of symmetries and the corresponding group of discrete symmetries.
Equation Miller’s Lie algebra surface $\Omega$ Lie algebra [De2] Discrete symmetries [De2]
${}_{2}F_{1}$ $sl(2,{\mathbb{C}})\oplus{\mathbb{C}}$ sphere $so(6,{\mathbb{C}})$ symmetries of cube
${}_{1}F_{1}$ $osc({\mathbb{C}})$ paraboloid $sch(2,{\mathbb{C}})$ ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$
${}_{2}F_{0}$ $sl(2,{\mathbb{C}})\oplus{\mathbb{C}}$ null quadric $sch(2,{\mathbb{C}})$ ${\mathbb{Z}}_{2}\times{\mathbb{Z}}_{2}$
${}_{0}F_{1}$ ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})\oplus{\mathbb{C}}$ paraboloid ${\mathbb{C}}^{2}\rtimes so(2,{\mathbb{C}})$ ${\mathbb{Z}}_{2}$
Gegenbauer $sl(2,{\mathbb{C}})\oplus{\mathbb{C}}$ sphere $so(5,{\mathbb{C}})$ symmetries of square
Hermite $osc({\mathbb{C}})$ cylinder $sch(1,{\mathbb{C}})$ ${\mathbb{Z}}_{4}$.
Above, $sch(n,{\mathbb{C}})$ denotes the complex Schrödinger Lie algebra, the Lie algebra of generalized symmetries of the heat equation in dimension $n$—see [De2].
The Lie algebra of the fourth column always contains the corresponding Miller’s Lie algebra (possibly, without the trivial term $\oplus{\mathbb{C}}$). Similarly, the group from the fifth column always contains the Weyl group of Miller’s Lie algebra.
7 Hypergeometric polynomials
Polynomial solutions of hypergeometric class equations will be called hypergeometric class polynomials. In this section we describe these solutions in detail.
We have already mentioned in (3.24) that there exist ladders of solutions of hypergeometric class equations consisting of polynomials. Note that according to our conventions, these ladders are descending. This is related to the usual convention for parameters of various types of hypergeometric functions: in order to get a polynomial of degree $n\in{\mathbb{N}}_{0}$, the parameter $a$ takes the value $-n$.
In Subsections 7.1 and 7.2 we describe the algebraic theory of hypergeometric class polynomials. They are centered around the so-called (generalized) Rodrigues formula.
[…]
7.1 2nd order differential operators with polynomial eigenfunctions
The following well-known and easy proposition shows that hypergeometric class operators have many polynomial eigenfunctions. In some sense, it is the property that characterizes this class among all second order differential operators.
Proposition 7.1
Let $\sigma(z)$, $\tau(z)$, $\mu(z)$ be arbitrary functions. Let $P_{n}(z)$, $n=0,1,\dots$, be polynomials such that $\deg P_{n}=n$ and $\eta_{n}\in{\mathbb{C}}$. Suppose that
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\mu(z)+\eta_{n}\right)P_{n}(z)=0,\ \ n=0,1,\dots$
Then $\sigma(z)$ is a polynomial of degree at most $2$, $\tau(z)$ is a polynomial of degree at most 1 and $\mu(z)$ a polynomial of degree 00 (a complex number). Thus, replacing $\mu+\eta_{n}$ with $\eta_{n}$, without limiting the generality we can assume that $\mu=0$. We then have
$\begin{array}[]{l}n(n-1)\frac{\sigma^{\prime\prime}}{2}+n\tau^{\prime}+\eta_{n}=0.\end{array}$ (7.1)
Here is another simple fact about polynomial solutions of hypergeometric class equations:
Proposition 7.2
Let $\deg\sigma\leq 2$ and $\deg\tau\leq 1$ and $\eta\in{\mathbb{C}}$ and
$k\frac{\sigma^{\prime\prime}}{2}+\tau^{\prime}\neq 0,\quad k=0,1,2,\dots.$ (7.2)
Then the space of polynomial solutions of
$\left(\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta\right)P(z)=0$ (7.3)
is at most $1$-dimensional.
Proof.  Suppose the space of polynomial solutions of (7.3) is 2-dimensional. We can assume that the degrees of these solutions are $0\leq n_{1}<n_{2}$. By (7.1),
$n_{i}(n_{i}-1)\frac{\sigma^{\prime\prime}}{2}+n_{i}\tau^{\prime}+\eta=0,\qquad i=1,2.$ (7.4)
Subtracting (7.4) for 2 and 1 and dividing by $n_{1}-n_{2}$ we obtain
$(n_{1}+n_{2}-1)\frac{\sigma^{\prime\prime}}{2}+\tau^{\prime}=0.$ (7.5)
Now possible values of $k:=n_{1}+n_{2}-1$ are $0,1,2,\dots$. This contradicts (7.2). $\Box$
7.2 Raising and lowering operators. Rodriguez formula
The following theorem shows how to construct hypergeometric class polynomials and describes some of their properties, which are easy to describe in a unified fashion. As usual, $\sigma$ denotes a polynomial of degree $\leq 2$, $\kappa$ a polynomial of degree $\leq 1$ and
1. and 2. of the following theorem is just a reformulation of Propositions 7.1 and 7.2, where instead of parameters $\tau,\eta$ we use $\kappa,\omega$.
Theorem 7.3
  1. 1.
    Suppose that $\omega\in{\mathbb{C}}$ and a polynomial of degree $n$ solves
    00 0 $=\Big{(}{\cal C}(\sigma,\kappa)+\omega\Big{)}v(z)$ (7.6)
    $=\Big{(}\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)+\kappa(z))\partial_{z}+\frac{\kappa^{\prime}}{2}+\omega\Big{)}v(z).$ (7.7)
    Then
    $n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\tfrac{1}{2})\kappa^{\prime}+\omega=0.$ (7.8)
  2. 2.
    Suppose that
    $k\frac{\sigma^{\prime\prime}}{2}+\kappa^{\prime}\neq 0,\quad k=2,3,\dots.$ (7.9)
    Then for any $\omega\in{\mathbb{C}}$ the space of polynomial solutions of (7.6) is at most $1$-dimensional.
  3. 3.
    Define
    $P_{n}(\sigma,\kappa;z)$ $:=\frac{1}{n!}\rho^{-1}(z)\partial_{z}^{n}\sigma^{n}(z)\rho(z)$ (7.10)
    $=\frac{1}{2\pi{\rm i}}\rho^{-1}(z)\int\limits_{[z^{+}]}\sigma^{n}(z+t)\rho(z+t)t^{-n-1}{\rm d}t,$ (7.11)
    where $[z^{+}]$ denotes a loop around $z$ in the counterclockwise direction. Then $P_{n}(\sigma,\kappa;z)$ is a polynomial of degree $n$ or less, and we have
    00 0 $=\Big{(}{\cal C}(\sigma,\kappa)-n(n+1)\frac{\sigma^{\prime\prime}}{2}-(n+\frac{1}{2})\kappa^{\prime}\Big{)}P_{n}(\sigma,\kappa;z)$ (7.12)
    $=\Big{(}\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)+\kappa(z))\partial_{z}-n(n+1)\frac{\sigma^{\prime\prime}}{2}-n\kappa^{\prime}\Big{)}P_{n}(\sigma,\kappa;z).$
Proof.  The integral representation (7.11) satisfies the assumptions of Theorem 5.1 about Euler transforms. This implies (7.12) holds. (7.10), usually called the Rodriguez formula, follows immediately from (7.11). $\Box$
Consider the descending ladder starting at $\kappa_{0}(z):=\kappa(z)$, $\omega_{0}=-\frac{\kappa^{\prime}}{2}$, that is
$\kappa_{-n}(z)$ $:=\kappa(z)-n\sigma^{\prime}(z),$ (7.13)
$\omega_{-n}$ $:=-(n+\tfrac{1}{2})\kappa^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}=-\kappa_{-n}^{\prime}(n+\tfrac{1}{2})-n(n+1)\frac{\sigma^{\prime\prime}}{2}.$ (7.14)
Proposition 7.4
The polynomials $P_{n}(\sigma,\kappa_{-n};z)$ satisfy the corresponding hypergeometric class equation
$\Big{(}{\cal C}(\sigma,\kappa_{-n})+\omega_{-n}\Big{)}P_{n}(\sigma,\kappa_{-n};z)$ $=0,$ (7.15)
the recurrence relations
$(\sigma(z)\partial_{z}+\kappa_{-n}(z))P_{n}(\sigma,\kappa_{-n};z)$ $=(n+1)P_{n+1}(\sigma,\kappa_{-n-1};z),$ (7.16)
$\partial_{z}P_{n+1}(\sigma,\kappa_{-n-1};z)$ $=\Big{(}n\frac{\sigma^{\prime\prime}}{2}+\kappa_{-n}^{\prime})p_{n}(\sigma,\kappa_{-n};z);$ (7.17)
and have a generating function
$\frac{\rho(z+t\sigma(z))}{\rho(z)}$ $=\sum\limits_{n=0}^{\infty}t^{n}P_{n}(\sigma,\kappa_{-n};z).$ (7.18)
7.3 Orthogonal polynomials
Consider $-\infty\leq a<b\leq+\infty$. Suppose that $]a,b[\ni x\mapsto\rho(x)$ is a positive measurable function. Define the Hilbert space
$L^{2}(]a,b[,\rho)$ :=$:=\Big{\{}f\quad\text{measurable on}\quad]a,b[\ |\ \int_{a}^{b}\rho(x)|f(x)|^{2}{\rm d}x<\infty\Big{\}},$ (7.19)
with the scalar product $(f|g):=\int_{a}^{b}\overline{f(x)}g(x)\rho(x){\rm d}x.$ (7.20)
Assume in addition
$\int_{a}^{b}\rho(x)|x|^{n}{\rm d}x<\infty,\quad n\in{\mathbb{N}}.$ (7.21)
Then the space of polynomials is contained in $L^{2}(]a,b[,\rho)$. Applying the Gram-Schmidt orthogonalization to the sequence $1,x,x^{2},\dots$ we can define an orthogonal family of polynomials $p_{0},p_{1},p_{2},\dots$ The following simple criterion is proven e.g. in [RS2]:
Theorem 7.5
Suppose that for some $\epsilon>0$
$\int_{a}^{b}{\rm e}^{\epsilon|x|}\rho(x){\rm d}x<\infty.$ (7.22)
Then polynomials are dense in $L^{2}(]a,b[,\rho)$. Therefore, the family $p_{0},p_{1},p_{2},\dots$ is an orthogonal basis of $L^{2}(]a,b[,\rho)$.
7.4 Hermiticity of Sturm-Liouville operators
In most of this paper we avoid using functional analysis. However, for hypergeometric class polynomials we will make an exception. In fact, often it is natural to view them as orthogonal bases consisting of eigenfunctions of certain self-adjoint operators.
Let us briefly recall some elements of the theory of operators on Hilbert spaces. Let ${\cal H}$ be a Hilbert space with the scalar product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). Let ${\cal A}$ be an operator on the domain ${\cal D}\subset{\cal H}$. We say that ${\cal A}$ is Hermitian on ${\cal D}$ in the sense of ${\cal H}$ if
$(f|{\cal A}g)=({\cal A}f|g),\quad f,g\in{\cal D}.$ (7.23)
(Unfortunately, in most of the contemporary mathematics literature, instead of Hermitian the word symmetric is used, which is a confusing misnomer). We say that ${\cal A}$ is self-adjoint if it is Hermitian and its spectrum is real. We say that it is essentially self-adjoint if it has a unique self-adjoint extension.
Let ${\cal A}$ be Hermitian. Suppose that $f_{i}\in{\cal D}$, $i=1,2$, and
${\cal A}f_{i}=\eta_{i}f_{i},\quad f_{i}\in{\cal D},\quad i=1,2.$ (7.24)
Then it is easy to see that $\eta_{i}$ are real, and if $\eta_{1}\neq\eta_{2}$, then $(f_{i}|f_{j})=0$. Therefore, eigenfunctions of Hermitian operators can be arranged into orthogonal families. If an operator is Hermitian and possesses an orthogonal basis of eigenvectors, then it is is essentially self-adjoint on linear combinations of its eigenvectors.
Second order differential operators on a segment of the real line are often called Sturm-Liouville operators. Let us make some remarks about the general theory of such operators, which will be useful in the analysis of orthogonality properties of hypegeometric class polynomials.
[…]
${\cal A}:=\sigma(x)\partial_{x}^{2}+\tau(x)\partial_{x}$ (7.25)
(Thus we consider an arbitrary Sturm-Liouville operator without the zeroth order term).
Let $\rho$ satisfy
$\sigma(x)\rho^{\prime}(x)=(\tau(x)-\sigma^{\prime}(x))\rho(x).$ (7.26)
Often it is convenient to rewrite (7.25) in the form
${\cal A}=\rho(x)^{-1}\partial_{x}\rho(x)\sigma(x)\partial_{x}.$ (7.27)
Suppose that $\sigma,\tau$ are real and $\rho$ is positive. It is useful to interpret ${\cal A}$ as an (unbounded) operator on the Hilbert space $L^{2}(]a,b[,\rho)$. Indeed, the following easy theorem shows that ${\cal A}$ is Hermitian on smooth functions that vanish close to the endpoints of the interval.
Theorem 7.6
Let
${\cal D}_{0}=\{f\in C^{\infty}(]a,b[)\ :\ f=0\ \ \hbox{in a neighborhood of}\ \ a,b\}.$ (7.28)
Then ${\cal A}$ is Hermitian on ${\cal D}_{0}$ in the sense of the Hilbert space $L^{2}(]a,b[,\rho)$.
7.5 Selecting endpoints for Sturm-Liouville operators
Unfortunately, the operator ${\cal A}$ is rarely essentially self-adjoint on (7.28). We will see that ${\cal A}$ is still Hermitian on functions that do not vanish near the endpoints, if these endpoints satisfy appropriate conditions:
Theorem 7.7
Let $-\infty<a<b<+\infty$ and
$\sigma(a)\rho(a)=\sigma(b)\rho(b)=0.$
Then ${\cal A}$ is Hermitian on the domain $C^{2}([a,b])$ in the sense of the space $L^{2}(]a,b[,\rho)$.
Proof.  Let $g,f\in C^{2}([a,b])$.
$(g|{\cal A}f)$ $=$ $\int_{a}^{b}\rho(x)\bar{g}(x)\rho(x)^{-1}\partial_{x}\sigma(x)\rho(x)\partial_{x}f(x){\rm d}x$
$=$ $\int_{a}^{b}\bar{g}(x)\partial_{x}\sigma(x)\rho(x)\partial_{x}f(x){\rm d}x$
$=$ $\bar{g(x)}\rho(x)\sigma(x)f^{\prime}(x)\Big{|}_{a}^{b}-\int_{a}^{b}(\partial_{x}\bar{g}(x))\sigma(x)\rho(x)\partial_{x}f(x){\rm d}x$
$=$ $-\bar{g^{\prime}(x)}\rho(x)\sigma(x)f(x)\Big{|}_{a}^{b}+\int_{a}^{b}(\partial_{x}\rho(x)\sigma(x)\partial_{x}\bar{g(x)})f(x){\rm d}x$
$=$ $\int_{a}^{b}\rho(x)\bar{\left(\rho(x)^{-1}\partial_{x}\sigma(x)\rho(x)\partial_{x}g(x)\right)}f(x){\rm d}x=({\cal A}g|f).$
$\Box$
Analogously we prove the following fact:
Theorem 7.8
Let
$\lim_{x\to-\infty}\sigma(x)\rho(x)|x|^{n}=\lim_{x\to+\infty}\sigma(x)\rho(x)|x|^{n}=0,\ \ \ n\in{\mathbb{N}}.$
Then ${\cal A}$ is Hermitian on the domain consisting of polynomial functions in the sense of the Hilbert space $L^{2}(]-\infty,\infty[,\rho)$.
Obviously, statements similar to Theorems 7.7 and 7.8 hold if $a=-\infty$ and $b$ is finite, or $a$ is finite and $b=\infty$.
We will see later that in applications to hypergeometric class equations the conditions of Theorems 7.7 and 7.8 are often sufficient to guarantee the essential self-adjointness of the operator ${\cal A}$ on the set of polynomials.
7.6 Orthogonality of hypergeometric polynomials
In some cases hypergeometric polynomials can be interpreted as an orthogonal basis of a certain weighted Hilbert space. In this case they are often called (very) classical orthogonal polynomials.
[…]
${\cal A}$ $:=\sigma(x)\partial_{x}^{2}+\tau(x)\partial_{x}=\rho(x)^{-1}\partial_{x}\rho(x)\sigma(x)\partial_{x}.$ (7.29)
is Hermitian on the space of polynomials in the sense of the Hilbert space $L^{2}(]a,b[,\rho)$. Therefore, polynomial eigenfunctions of ${\cal A}$ are pairwise orthogonal (at least if the corresponding eigenvalues are distinct, which is generically the case, see (7.8)). The following theorem says more: we compute the square norm of the polynomials obtained from the Rodriguez formula.
Theorem 7.9
$\int_{a}^{b}P_{m}(\sigma,\kappa;x)P_{n}(\sigma,\kappa;x)\rho(x){\rm d}x=\frac{\delta_{mn}}{n!}$ $\prod_{j=1}^{n}\Big{(}-\kappa^{\prime}+j\frac{\sigma^{\prime\prime}}{2}\Big{)}\int_{a}^{b}\sigma^{n}(x)\rho(x){\rm d}x.$ (7.30)
Proof.  It is enough to assume that $m\leq n$. Write $P_{n}(x)$ for $P_{n}(\sigma,\kappa;x)$.
$\int_{a}^{b}P_{m}(x)P_{n}(x)\rho(x){\rm d}x$ (7.31)
$=$ $\frac{1}{n!}\int_{a}^{b}P_{m}(x)\partial_{x}^{n}\sigma^{n}(x)\rho(x){\rm d}x$ (7.32)
$=$ $\frac{(-1)^{n}}{n!}\int_{a}^{b}\Big{(}\partial_{x}^{n}P_{m}(x)\Big{)}\sigma^{n}(x)\rho(x){\rm d}x.$ (7.33)
This is zero if $m<n$ and for $m=n$ we use
$\partial_{x}^{n}P_{n}(x)=\prod_{j=1}^{n}\Big{(}\kappa^{\prime}-j\frac{\sigma^{\prime\prime}}{2}\Big{)}.$ (7.34)
$\Box$
7.7 Review of types of hypergeometric class polynomials
In the remaining part of the section we review various types of hypergeometric polynomials. We will discuss only the properties that follow directly from the general theory described in previous subsections. Here is the list of items that we will cover:
  • The choice of $\sigma,\kappa$ and the corresponding weight $\rho$.
  • The Rodriguez formula (7.10) defining the polynomial and its expression in terms of hypergeometric class functions.
  • The degree of the polynomial, see (7.1).
  • The differential equation, see (7.12).
  • The pair of recurrence relations that follow from the Rodriguez formula, see (7.16) and (7.17).
  • The generating function related to the Rodriguez formula, see (7.18).
  • The choice of endpoints for which the corresponding Sturm-Liouville operator is essentially self-adjoint on the space of polynomials (if applicable), see Theorems 7.7 , 7.8 and 7.5.
  • The square norm (if applicable), see Theorem 7.9.
Let us mention that hypergeometric class polynomials possess various useful features that will not be listed in the following subsections. For instance, they typically have additional recurrence relations and generating functions. We list only those that follow directly from the general theory described above.
7.8 Jacobi polynomials
Consider $\alpha,\beta\in{\mathbb{C}}$ and
$\sigma(z)=1-z^{2},\quad\kappa(z)=\alpha(1-z)+\beta(1+z),\quad\rho(z)=(1-z)^{\alpha}(1+z)^{\beta}.$ (7.35)
For $n\in\{0,1,\dots\}$ set
$P_{n}^{\alpha,\beta}(x)$ $=$ $\frac{(-1)^{n}}{2^{n}n!}(1-x)^{-\alpha}(1+x)^{-\beta}\partial_{x}^{n}(1-x)^{\alpha+n}(1+x)^{\beta+n}$ (7.36)
$=$ $\frac{(1+\alpha)_{n}}{n!}{}_{2}F_{1}\Big{(}-n,n+\alpha+\beta+1;\alpha+1;\frac{1-x}{2}\Big{)}.$ (7.37)
Then $P_{n}^{\alpha,\beta}$ is a polynomial of degree at most $n$. More precisely:
  1. 1.
    If $\alpha+\beta\not\in\{-2n,\dots,-n-1\}$, then $\deg P_{n}^{\alpha,\beta}=n$. It is then up to a coefficient the unique eigensolution of the operator
    ${\cal A}:=(1-x^{2})\partial_{x}^{2}+(\beta-\alpha-(\alpha+\beta+2)x)\partial_{x},$ (7.38)
    which is a polynomial of degree $n$.
  2. 2.
    If $\alpha+\beta\in\{-2n,\dots,-n-1\}$, but $\alpha\not\in\{-n,\dots,-1\}$ (or, equivalently, $\beta\not\in\{-n,\dots,-1\}$), then $\deg P_{n}^{\alpha,\beta}=-\alpha-\beta-n-1$.
  3. 3.
    If $\alpha+\beta\in\{-2n,\dots,-n-1\}$, but $\alpha\in\{-n,\dots,-1\}$ (or, equivalently, $\beta\in\{-n,\dots,-1\}$), then $P_{n}^{\alpha,\beta}=0$.
$P_{n}^{\alpha,\beta}$ satisfy the Jacobi equation, which is a slightly modified ${}_{2}F_{1}$ equation
$\left((1-x^{2})\partial_{x}^{2}+(\beta-\alpha-(\alpha+\beta+2)x)\partial_{x}+n(n+\alpha+\beta+1)\right)P_{n}^{\alpha,\beta}(x)=0.$
Recurrence relations:
$\partial_{x}P_{n}^{\alpha,\beta}(x)$ $=$ $\frac{\alpha+\beta+n+1}{2}P_{n-1}^{\alpha+1,\beta+1},$ (7.39)
$-\frac{(1-x^{2})\partial_{x}+\beta-\alpha-(\alpha+\beta)x}{2}P_{n}^{\alpha,\beta}(x)$ $=$ $(n+1)P_{n+1}^{\alpha-1,\beta-1}(x).$ (7.40)
Generating function:
$\sum_{n=0}^{\infty}P_{n}^{\alpha-n,\beta-n}(x)2^{n}t^{n}=\bigl{(}1+t(1+x)\bigr{)}^{\alpha}\bigl{(}1-t(1-x)\bigr{)}^{\beta}.$ (7.41)
If $\alpha,\beta>-1$, then ${\cal A}$ is self-adjoint on the space of polynomials in the sense of
$L^{2}(]-1,1[,(1-x)^{\alpha}(1+x)^{\beta}).$
The Jacobi polynomials are its eigenfunctions and form an orthogonal basis with the normalization
$\int_{-1}^{1}(P_{n}^{\alpha,\beta}(x))^{2}(1-x)^{\alpha}(1+x)^{\beta}{\rm d}x=\frac{\Gamma(1+\alpha+n)\Gamma(1+\beta+n)2^{\alpha+\beta+1}}{(1+2n+\alpha+\beta)n!\Gamma(1+\alpha+\beta+n)}.$ (7.42)
7.9 Laguerre polynomials
Consider $\alpha\in{\mathbb{C}}$ and
$\sigma(z)=z,\quad\kappa(z)=\alpha-z,\quad\rho(z)={\rm e}^{-z}z^{\alpha}.$ (7.43)
For $n\in{\mathbb{N}}$ set
$L_{n}^{\alpha}(x)$ $=$ $\frac{1}{n!}{\rm e}^{x}x^{-\alpha}\partial_{x}^{n}{\rm e}^{-x}x^{n+\alpha}$
$=$ $\frac{(1+\alpha)_{n}}{n!}{}_{1}F_{1}(-n;1+\alpha;x).$
Then $L_{n}^{\alpha}$ is a polynomial of degree $n$. It is a unique (up to a coefficient) eigenfunction of the operator
${\cal A}:=x\partial_{x}^{2}+(\alpha+1-x)\partial_{x},$ (7.44)
which is a polynomial of degree $n$. $L_{n}^{\alpha}$ satisfy the Laguerre equation, which is the ${}_{1}F_{1}$ equation with modified parameters:
$\left(x\partial_{x}^{2}+(\alpha+1-x)\partial_{x}+n\right)L_{n}^{\alpha}(x)=0.$
Recurrence relations:
$(x\partial_{x}+\alpha-x)L_{n}^{\alpha}(x)$ $=$ $(n+1)L_{n+1}^{\alpha-1}(x),$ (7.45)
$\partial_{x}L_{n}^{\alpha}(x)$ $=$ $-L_{n-1}^{\alpha+1}(x).$ (7.46)
Generating function:
${\rm e}^{-tz}(1+t)^{\alpha}=\sum\limits_{n=0}^{\infty}t^{n}L_{n}^{\alpha-n}(z).$ (7.47)
If $\alpha>-1$, then ${\cal A}$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}([0,\infty[,{\rm e}^{-x}x^{\alpha})$. The Laguerre polynomials are its eigenfunctions and form an orthonormal basis with the normalization
$\int_{0}^{\infty}L_{n}^{\alpha}(x)^{2}x^{\alpha}{\rm e}^{-x}{\rm d}x=\frac{\Gamma(1+\alpha+n)}{n!}.$
7.10 Bessel polynomials
Consider $\theta\in{\mathbb{C}}$ and
$\sigma(z)=z^{2},\quad\kappa(z)=-1+\theta z,\quad\rho(z)={\rm e}^{-z^{-1}}z^{\theta}.$ (7.48)
For $n=0,1,\dots$ set
$B_{n}^{\theta}(z):$ $=\frac{1}{n!}z^{-\theta}{\rm e}^{z^{-1}}\partial_{z}^{n}{\rm e}^{-z^{-1}}z^{\theta+2n}$ (7.49)
$=\frac{1}{n!}{}_{2}F_{0}(-n,n+\theta+1;-;z)$ (7.50)
$=(-z)^{n}L_{n}^{-\theta-2n-1}(-z^{-1}).$ (7.51)
Then $B_{n}^{\theta}$ is a polynomial of degree $n$. It is a unique (up to a coefficient) eigenfunction of the operator
${\cal A}:=z^{2}\partial_{z}^{2}+(-1+(2+\theta)z)\partial_{z},$ (7.52)
which is a polynomial of degree $n$. $B_{n}^{\theta}$ satisfy the ${}_{2}F_{0}$ equation with adjusted parametes:
$\Big{(}z^{2}\partial_{z}^{2}+(-1+(2+\theta)z)\partial_{z}-\frac{1}{2}n(1+\theta+n)\Big{)}B_{n}^{\theta}(z)$ $=$ $0$
Recurrence relations:
$\partial_{z}B_{n}^{\theta}(z)$ $=$ $-(n+\theta+1)B_{n-1}^{\theta+2}(z),$
$\left(z^{2}\partial_{z}-1-\theta z\right)B_{n}^{\theta}(z)$ $=$ $-(n+1)B_{n+1}^{\theta-2}(z).$
Generating function:
$(1+tz)^{\theta}\exp\Big{(}\frac{-t}{1+tz}\Big{)}=\sum\limits_{n=0}^{\infty}t^{n}B_{n}^{\theta-2n}(z).$ (7.53)
Bessel polynomials do not form an orthogonal basis on any interval.
7.11 Hermite polynomials
Consider
$\sigma(z)=1,\quad\kappa(z)=-2z,\quad\rho(z)={\rm e}^{-z^{2}}.$ (7.54)
For $n=0,1,\dots$ set
$H_{n}(x)$ $=\frac{(-1)^{n}}{n!}{\rm e}^{x^{2}}\partial_{x}^{n}{\rm e}^{-x^{2}}$ (7.55)
$=\frac{2^{n}}{n!}S(-n;x),$ (7.56)
where $S(a,x)$ is the Hermite function defined in (6.32). Then $H_{n}$ is a polynomial of degree $n$ and is (up to a multiplicative constant) the only eigenfunction of the operator
${\cal A}:=\partial_{x}^{2}-2x\partial_{x},$ (7.57)
which is a polynomial of degree $n$. It satisfies the Hermite equation
$(\partial_{x}^{2}-2x\partial_{x}+2n)H_{n}(x)=0.$
Recurrence relations:
$(-\partial_{x}+2x)H_{n}(x)$ $=(n+1)H_{n+1}(x),$ (7.58)
$\partial_{x}H_{n}(x)$ $=2H_{n-1}(x).$ (7.59)
Generating function:
$\sum_{n=0}^{\infty}t^{n}H_{n}(x)$ $={\rm e}^{2tx-t^{2}}.$ (7.60)
The operator ${\cal A}$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}({\mathbb{R}},{\rm e}^{-x^{2}})$. The Hermite polynomials are its eigenfunctions and form an orthogonal basis with the normalization
$\int_{-\infty}^{\infty}H_{n}(x)^{2}{\rm e}^{-x^{2}}{\rm d}x=\frac{\sqrt{\pi}2^{n}}{n!}.$
Remark 7.10
The definition of Hermite polynomials that we gave is consistent with the generalized Rodrigues formula (7.10). In the literature one can also find other conventions for Hermite polynomials, e.g. $H_{n}(x):=(-1)^{n}{\rm e}^{x^{2}}\partial_{x}^{n}{\rm e}^{-x^{2}}.$
8 Degenerate case
If two indices of a regular-singular (also called Fuchsian) point of a differential equation differ by an integer, then the usual Frobenius method [WW] produces in general only one solution, up to a coefficient. We call this case degenerate. In this section we will be discuss the degenerate case of hypergeometric class equations. Without limiting the generality, the regular-singular point will be 00.
8.1 United hypergeometric function in the degenerate case
Suppose that we are in the setting of Section 4.3. That means, $\sigma(0)=0$, $\sigma^{\prime}(0)=1$, and we set $\kappa(0)=m$. Thus,
$\sigma(z)=\frac{\sigma^{\prime\prime}}{2}z^{2}+z,\quad\kappa(z)=m+\kappa^{\prime}z,\quad\omega$ (8.1)
are the parameters of the equation we consider, see (4.7). As in (4.11), we also introduce the transformed parameters:
$\kappa^{\circ}(z):=-m+(\kappa^{\prime}-m\sigma^{\prime\prime})z,\quad\omega^{\circ}:=\omega-m\kappa^{\prime}+m^{2}\frac{\sigma^{\prime\prime}}{2},\quad m^{\circ}=-m.$ (8.2)
We assume in addition that $m$ is an integer. Using
$\frac{1}{\Gamma(m+1)}=\begin{cases}0&m=\dots,-2,-1;\\ \frac{1}{m!},&m=0,1,\dots,\end{cases}$ (8.3)
we obtain
${\mathbf{F}}(\sigma,\kappa,\omega;z)=\sum_{n=\max\{0,-m\}}^{\infty}\frac{\Pi_{j=0}^{n-1}(\omega+(j+\frac{1}{2})\kappa^{\prime}+\frac{j(j+1)}{2}\sigma^{\prime\prime})}{(n+m)!n!}(-z)^{n}.$ (8.4)
This easily implies the identity
${\mathbf{F}}(\sigma,\kappa^{\circ},\omega^{\circ};z)=$ $\prod_{j=0}^{m-1}\Big{(}\omega-\kappa^{\prime}(j+\tfrac{1}{2})+\frac{\sigma^{\prime\prime}}{2}j(j+1)\Big{)}(-z)^{m}{\mathbf{F}}(\sigma,\kappa,\omega;z),\quad m\in{\mathbb{N}}_{0}.$ (8.5)
Therefore, for integer $m$ the functions $\mathbf{F}(\sigma,\kappa,\omega;z)$ and $z^{-m}{\mathbf{F}}(\sigma,\kappa^{\circ},\omega^{\circ};z)$ are proportional to one another and do not form a basis of solutions.
8.2 The power-exponential function
In order to describe the theory of the degenerate case in a unified way, it will be convenient first to unify the power and exponential function in a single function. More precisely, let $a,\mu\in{\mathbb{C}}$. Consider the function
$f(a,\mu;u):=\left\{\begin{array}[]{ll}(1+\mu u)^{a/\mu},&\ \ \mu\neq 0,\\[8.53581pt] {\rm e}^{au},&\ \ \mu=0.\end{array}\right.$
It is analytic for $u\neq-\mu^{-1}$. From now on we will write $(1+\mu u)^{a/\mu}$ instead of $f(a,\mu;u)$.
Let us list some properties of this function.
Theorem 8.1
(1) $(1+\mu u)^{a/\mu}$ is the unique solution of
$\big{(}(1+\mu u)\partial_{u}-a\big{)}f(u)=0,$ (8.6)
$f(0)=1.$ (8.7)
(2) For $|u|<|\mu|^{-1}$ we have
$(1+\mu u)^{a/\mu}=\sum_{n=0}^{\infty}\frac{a(a-\mu)\cdots(a-(n-1)\mu)}{n!}u^{n},$
$(1+\mu u)^{a_{1}/\mu}(1+\mu u)^{a_{2}/\mu}=(1+\mu u)^{(a_{1}+a_{2})/\mu}.$
8.3 Generating functions
Let us now fix $\sigma(z)=\frac{\sigma^{\prime\prime}}{2}z^{2}+z$, as in (8.1). Fix also $\delta,\theta$. Let $\rho$ and $\tilde{\rho}$ solve the equations
$-\frac{\sigma^{\prime\prime}}{2}\delta=\rho\tilde{\rho},\qquad\rho+\tilde{\rho}=\theta.$ (8.8)
Consider two families of hypergeometric class operators
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\theta\Big{)}z\Big{)}\partial_{z}-m\rho-\delta,$ (8.9)
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\theta\Big{)}z\Big{)}\partial_{z}-m\tilde{\rho}-\delta.$ (8.10)
In terms of our standard parameters, these operators can be written as
${\mathcal{C}}(\sigma,\kappa_{m})+\omega_{m},$ (8.11)
${\mathcal{C}}(\sigma,\kappa_{m})+\tilde{\omega}_{m},$ (8.12)
where
$\kappa_{m}(z)$ $:=m+\Big{(}\frac{\sigma^{\prime\prime}}{2}(m-1)-\theta\Big{)}z,$ (8.13)
$\omega_{m}$ $:=\frac{\theta}{2}-\delta-m\rho-\frac{\sigma^{\prime\prime}}{4}(m-1),$ (8.14)
$\tilde{\omega}_{m}$ $:=\frac{\theta}{2}-\delta-m\tilde{\rho}-\frac{\sigma^{\prime\prime}}{4}(m-1).$ (8.15)
Note that
$\kappa_{m}^{\circ}=\kappa_{-m},\quad\omega_{m}^{\circ}=\tilde{\omega}_{-m},\quad\tilde{\omega}_{m}^{\circ}=\omega_{-m}.$ (8.16)
We are going to construct solutions to the equations given by (8.9) and (8.10). In order to present them in a unified way, it is convenient to choose $a,b,\mu,\nu$ satisfying
$\rho=\mu b,\quad\tilde{\rho}=\nu a,\quad\delta=ab.$ (8.17)
By $[(\alpha,\beta)^{+}]$ we will denote the counterclockwise loop around $\alpha,\beta\in{\mathbb{C}}$. We define
$\Psi_{m}(z)$ $:=\frac{1}{2\pi{\rm i}}\int_{[(0,-z\nu))^{+}]}(1+\mu u)^{-\frac{a}{\mu}}\Big{(}1+\frac{\nu z}{u}\Big{)}^{-\frac{b}{\nu}}u^{-m-1}{\rm d}u,$ (8.18)
$\tilde{\Psi}_{m}(z)$ $:=\frac{1}{2\pi{\rm i}}\int_{[(0,-z\mu)^{+}]}\Big{(}1+\frac{\mu z}{v}\Big{)}^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}v^{-m-1}{\rm d}v.$ (8.19)
Theorem 8.2
  1. 1.
    For $|z\nu|<|u|<|\mu^{-1}|$ we have
    $(1+\mu u)^{-\frac{a}{\mu}}\Big{(}1+\frac{\nu z}{u}\Big{)}^{-\frac{b}{\nu}}=\sum_{m\in{\mathbb{Z}}}u^{m}\Psi_{m}(z).$ (8.20)
    and for $|z\mu|<|u|<|\nu^{-1}|$ we have
    $\Big{(}1+\frac{\mu z}{v}\Big{)}^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}=\sum_{m\in{\mathbb{Z}}}v^{m}\tilde{\Psi}_{m}(z).$ (8.21)
  2. 2.
    $\Psi_{m}(z)=z^{-m}\tilde{\Psi}_{-m}(z).$ (8.22)
  3. 3.
    For $m\geq 0$,
    $\Psi_{m}(0)$ $=(-1)^{m}\frac{a(a+\mu)\cdots(a+\mu(m-1))}{m!},$ (8.23)
    $\tilde{\Psi}_{m}(0)$ $=(-1)^{m}\frac{b(b+\nu)\cdots(b+(m-1)\nu)}{m!}.$ (8.24)
  4. 4.
    The functions $\Psi_{m}(z)$ and $\tilde{\Psi}_{m}(z)$ satisfy the following hypergeometric class differential equations:
    $\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu(1+m)+a\nu+b\mu)z)\partial_{z}-(m\mu b+ab)\right)\Psi_{m}(z)$ $=0,$ (8.25)
    $\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu(1+m)+a\nu+b\mu)z)\partial_{z}-(m\nu a+ab)\right)\tilde{\Psi}_{m}(z)$ $=0.$ (8.26)
  5. 5.
    The functions $\Psi_{m}(z)$ and $\tilde{\Psi}_{m}(z)$ are proportional to hypergeometric functions:
    $\Psi_{m}(z)$ $=(-1)^{m}a(a+\mu)\cdots(a+\mu(m-1))\mathbf{F}(\sigma,\kappa_{m},\omega_{m};z);$ (8.27)
    $\tilde{\Psi}_{m}(z)$ $=(-1)^{m}b(b+\nu)\cdots(b+(m-1)\nu)\mathbf{F}(\sigma,\kappa_{m},\tilde{\omega}_{m};z);$ (8.28)
Proof.  0. follows immediately from the definition and the Laurent expansion.
We can rewrite (8.20) as
$(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}=\sum_{m\in{\mathbb{Z}}}u^{m}\Psi_{m}(uv).$ (8.29)
By (8.21) we have
$(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}=\sum_{m\in{\mathbb{Z}}}v^{m}\tilde{\Psi}_{m}(uv)=\sum_{m\in{\mathbb{Z}}}u^{-m}(uv)^{m}\tilde{\Psi}_{m}(uv).$ (8.30)
This implies 1. 2. is immediate.
Let us show 3.. We have the identity
$\begin{array}[]{rl}0&=\Big{(}\partial_{u}((1+\nu v)\partial_{v}+b)+\partial_{v}((1+\mu u)\partial_{u}+a)\\[8.53581pt] &-((1+\mu u)\partial_{u}+u)((1+\nu v)\partial_{v}+b)\Big{)}(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}\\[8.53581pt] &=\left(\partial_{u}\partial_{v}-\mu\nu uv\partial_{u}\partial_{v}-\mu bu\partial_{u}-a\nu v\partial_{v}-ab\right)(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}.\end{array}$ (8.31)
Besides, if $z=uv$, then
$u\partial_{u}u^{m}\Psi_{m}(uv)=u^{m}(m+z\partial_{z})\Psi_{m}(z),$
$v\partial_{v}u^{m}\Psi_{m}(uv)=u^{m}z\partial_{z}\Psi_{m}(z),$
$\partial_{v}\partial_{u}u^{m}\Psi_{m}(uv)=u^{m}\left((m+1)\partial_{z}+z\partial_{z}^{2}\right)\Psi_{m}(z),$
Hence
$\begin{array}[]{l}\left(\partial_{u}\partial_{v}-\mu\nu uv\partial_{u}\partial_{v}-\mu bu\partial_{u}-a\nu v\partial_{v}-ab\right)u^{m}\Psi_{m}(uv)\\[5.69054pt] =u^{m}\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu+(m\mu+a)\nu+b\mu)z)\partial_{z}-(m\mu+a)b\right)\Psi_{m}(z).\end{array}$
Thus
$0=\sum_{m\in{\mathbb{Z}}}u^{m}\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu+(m\mu+a)\nu+b\mu)z)\partial_{z}-(m\mu+a)b\right)\Psi_{m}(z),$
which implies (8.25). $\Box$
In the remaining part of the section we describe separately three types of degenerate hypergeometric class functions.
8.4 The ${}_{2}F_{1}$ function
For $m\in{\mathbb{Z}}$ we have
${\bf F}(a,b;1+m;z)$ $=\sum_{n=\max(0,-m)}\frac{(a)_{n}(b)_{n}}{n!(m+n)!}z^{n},$ (8.32)
$(a-m)_{m}(b-m)_{m}{\bf F}(a,b;1+m;z)$ $=z^{-m}{\bf F}(a-m,b-m;1-m;z).$ (8.33)
We have an integral representation and a generating function:
$\frac{1}{2\pi{\rm i}}\int\limits_{[(0,z)^{+}]}(1-t)^{-a}\Big{(}1-\frac{z}{t}\Big{)}^{-b}t^{-m-1}{\rm d}t$ $=(a)_{m}{\bf F}(a+m,b;1+m;z),$ (8.34)
$(1-t)^{-a}\Big{(}1-\frac{z}{t}\Big{)}^{-b}$ $=\sum_{m\in{\mathbb{Z}}}t^{m}(a)_{m}{\bf F}(a+m,b;1+m).$ (8.35)
8.5 The ${}_{1}F_{1}$ function
If $m\in{\mathbb{Z}}$, we have
${\bf F}(a;1+m;z)$ $=\sum_{n=\max(0,-m)}\frac{(a)_{n}}{n!(m+n)!}z^{n},$ (8.36)
$(a-m)_{m}{\bf F}(a;1+m;z)$ $=z^{-m}{\bf F}(a-m;1-m;z).$ (8.37)
We have two integral representations and the corresponding generating functions:
$\frac{1}{2\pi{\rm i}}\int\limits_{[(z,0)^{+}]}{\rm e}^{t}\Big{(}1-\frac{z}{t}\Big{)}^{-a}t^{-m-1}{\rm d}t$ $=$ ${\bf F}(a,1+m;z),$
$\frac{1}{2\pi{\rm i}}\int\limits_{[(0,1)^{+}]}{\rm e}^{\frac{z}{t}}(1-t)^{-a}t^{-m-1}{\rm d}t$ $=$ $z^{-m}{\bf F}(a,1-m;z),$
${\rm e}^{t}\Big{(}1-\frac{z}{t}\Big{)}^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{\bf F}(a;m;z)),$
${\rm e}^{\frac{z}{t}}(1-t)^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}z^{-m}{\bf F}(a;1-m;z).$
8.6 The ${}_{0}F_{1}$ function
If $m\in{\mathbb{Z}}$, then
${\bf F}(1+m;z)$ $=\sum_{n=\max(0,-m)}\frac{1}{n!(m+n)!}z^{n},$ (8.38)
${\bf F}(1+m;z)$ $=z^{-m}{\bf F}(1-m;z).$ (8.39)
We have an integral representation, called the Bessel formula, and a generating function:
$\frac{1}{2\pi{\rm i}}\int\limits_{[0^{+}]}{\rm e}^{t+\frac{z}{t}}t^{-m-1}{\rm d}t$ $={\bf F}(1+m;z),$ (8.40)
${\rm e}^{t}{\rm e}^{\frac{z}{t}}$ $=\sum_{m\in{\mathbb{Z}}}t^{m}{\bf F}(1+m;z).$ (8.41)
Acknowledgement. This work was supported by National Science Center (Poland) under the grant UMO-2019/35/B/ST1/01651. I thank Rupert Frank for an inspiring remark. I am also grateful to Adam Kaliński for useful comments on the manuscript.